### v1

# Online Makespan Scheduling under Scenarios

Abstract
We consider a natural extension of online makespan scheduling on identical parallel machines by introducing scenarios. A scenario is a subset of jobs, and the task of our problem is to find a global assignment of the jobs to machines so that the maximum makespan under a scenario, i.e., the maximum makespan of any schedule restricted to a scenario, is minimized.
For varying values of the number of scenarios and machines, we explore the competitiveness of online algorithms. We prove tight and near-tight bounds, several of which are achieved through novel constructions. In particular, we leverage the interplay between the unit processing time case of our problem and the hypergraph coloring problem both ways: We use hypergraph coloring techniques to steer an adversarial family of instances proving lower bounds for our problem, which in turn leads to lower bounds for several variants of online hypergraph coloring.
keywords:
online scheduling, scenario-based model, online algorithms
category:
[…]
1 Introduction
We study a natural extension of online makespan scheduling, also known as online load balancing, one of the most well-known problems in the scope of online optimization. In online makespan scheduling, we know a number $m$ of identical parallel machines in advance, in addition to which jobs $j$ are revealed one at a time along with their processing times $p_{j}\geq 0$. Our task is to assign the jobs to a machine as soon as they are revealed, solely under the knowledge of the preceding jobs, so as to minimize the makespan, i.e., the maximum sum of processing times any machine is assigned to. This problem was initially studied by Graham almost 60 years ago, who proved that assigning every job to a currently least loaded machine is a $(2-\frac{1}{m})$-competitive algorithm [graham]. It was observed that no better deterministic algorithm can exist for $m\in\{2,3\}$ [graham][grahamlb], while for larger values of $m$, such a tight ratio is still not known.
We extend online makespan scheduling by a scenario-based model: We are given a number $m$ of machines. There is a known number $K$ of job sets (called scenarios) $S_{k}\subseteq\{1,\ldots,n\}$, $k\in[K]$, over a common ground set $\{1,\ldots,n\}$. Although the number of scenarios is known, neither the number of jobs nor their processing times $p_{j}$ ($j\in[n]$) are known in advance. The ground set $[n]$ is revealed in increasing order of jobs; more precisely, we are incrementally informed of $p_{1},\ldots,p_{j}$ as well as $S_{k}\cap\{1,\ldots,j\}$ for increasing $j$. The task is to find a global assignment $\tau\colon\{1,\ldots,n\}\to\{1,\ldots,m\}$ of the jobs to machines that delivers a reasonable solution restricted to every scenario. To be more precise, the objective is to minimize the makespan of the worst-case scenario, i.e., the scenario that yields the highest makespan. As in the classical online scheduling problem, a job $j$ must be assigned irrevocably to a machine $\tau(j)$ as soon as it is revealed. In essence, we are solving a finite number of instances of Online Makespan Scheduling over the same ground set of jobs simultaneously.
The simplicity of the scenario model offers a variety of interpretations and applications. For instance, we can view the scenarios as distinct services that are provided in $m$ (number of machines) sites to $n$ (number of jobs) customers, where the customers are revealed one by one with their desired services and need to be assigned to one of the sites immediately. Another interpretation of our problem may concern a fair balancing of goods according to multiple agents in the following setting: Each of the $K$ agents are interested in partitioning a subset $S_{k}$, $k\in\{1,\ldots,K\}$ of $n$ goods with weights $p_{j}$ among $m$ buckets, where the partition should be as balanced for all agents as possible (i.e., we would like to minimize the maximum bucket load for any agent). Here, the goods are revealed one by one together with the agents that are interested in them.
Moreover, even special cases of online makespan scheduling under scenarios correspond to interesting questions in their own regard. The special case $p_{j}\equiv 1$ of unit processing times is related to some cases of discrepancy minimization as well as hypergraph coloring.
In the large array of previous research on combinatorial optimization under scenarios, it was shown for several offline problems that optimizing under a number of scenarios is computationally harder than their underlying single-scenario counterparts. Accordingly, our goal in this work is to gain a better understanding of the intersection of scenario-based models and competitiveness, in particular how competitiveness diminishes under a growing number of scenarios.
1.1 Our Contribution
Implementing a miscellany of ideas to analyze competitiveness bounds for the varying numbers $m$ of machines and $K$ of scenarios, we achieve several tight bounds and further nontrivial bounds. An overview of our results, depending on the numbers $m$ and $K$, can be seen in Table 1. Several of the entries are obtained via a generalization of Graham’s List Scheduling Algorithm which we extend using the pigeonhole principle.
By a simple example, one can show that for any number $m\geq 2$ of machines and (at least) $3$ scenarios, no algorithm with a competitive ratio better than $2$ can exist, even for jobs with unit processing times. To underline the competitiveness gap between two and three scenarios, we contrast the aforementioned result with an algorithm for two machines and two scenarios, which is one of our main results.
Theorem 1.1.
There exists a $\nicefrac{{5}}{{3}}$-competitive algorithm for Online Makespan Scheduling under Scenarios for $m=K=2$.
This algorithm as well as its analysis combine techniques well-known from algorithms for online makespan scheduling with novel ideas that concern the scenario-based properties of a schedule. We first alter the definition of competitive ratio by allowing only specific lower bounds for an offlie optimum, which allows us to make structural modifications and restrictions on worst-case instances. What sets our analysis apart from previous work is that we alternate between forward and reverse engineering: Assuming that our algorithm obeys some fixed rules, we explore what we may assume about a well-structured worst-case instance. In spite of seemingly strong restrictions that we impose, the ratio of our algorithm is not far from the best achievable bound by any deterministic algorithm, which we show is at least $\frac{9+\sqrt{17}}{8}\approx 1.640$. Altogether, we obtain a near-complete picture of competitiveness for $m=2$.
The remainder of the paper demonstrates a disparity between increasing the number $m$ of machines and increasing the number $K$ of scenarios. For a fixed but arbitrary $m$, we show that there exists a sufficiently large number $K(m)$ of scenarios and a family of instances with $K(m)$ scenarios, on which no deterministic online algorithm can achieve a ratio better than $m$. Furthermore, we can even choose all instances to have unit processing times $p_{j}\equiv 1$. This result is particularly noteworthy, considering that any online algorithm on $m$ machines is trivially $m$-competitive: Indeed, the online output is at most the total load and an offline optimum is at least the average load, and these are by at most a factor of $m$ apart.
To find the suitable number $K(m)$, we interpret the instances as hypergraphs, passing to a problem we call Online Makespan Hypergraph Coloring$(m)$ instead (cf. Section 1.2). Strikingly, such a family can be created using hypertrees on which no deterministic algorithm can achieve a coloring without a monochromatic edge simultaneously.
Theorem 1.2.
Let $m\in\mathbb{N}$. There exists a number $K(m)$ such that for any $r<m$, no deterministic algorithm for Online Makespan Hypergraph Coloring$(m)$ is $r$-competitive, even when restricted to hypertrees with $K(m)$ hyperedges.
The statement of Theorem 1.2 is best possible in the sense that the numbers $K(m)$ cannot be replaced by a global $K$ for which the incompetitiveness holds for all $m$ (cf. Theorem 2.3).
The proof of Theorem 1.2 can be found in the full version of the paper. To showcase some of our techniques, we sketch a proof for a similar statement for $K=3$ in Section 4. We utilize a subhypergraph on $7$ nodes as a gadget and obtain a relatively compact construction, both in terms of the hypergraph and the case distinction that we apply to analyze it.
[…]
Corollary 1.
Let $n\in\mathbb{N}$. There exists no $o(\mathrm{slog_{5}}n)$-competitive deterministic algorithm for Online Hypergraph Coloring with nodes $v_{1},\ldots,v_{n}$ (revealed in this order) even if we restrict to hypertrees and additionally, the subsets $e\cap\{v_{1},\ldots,v_{j}\}$ of hyperedges $e\in E$ are known upon the assignment of node $v_{j}$.
For unit processing times, $K=3$ and varying $m$, we further propose a $2$-competitive algorithm, which is best possible due to the lower bounds mentioned earlier. Although the special case of unit processing times is admittedly a restriction, the result still illustrates a contrast with the tight lower bound of $m$ for a varying number of scenarios, which also holds already for instances with unit processing times.
m $\downarrow$ K $\rightarrow$ 2 3 4+
2 $>1.640$, $\leq\nicefrac{{5}}{{3}}$ $=2$ $=2$
3 $\leq 2$ $\geq 2$, $\geq 2$
$=2$ for $p_{j}\equiv 1$ $=3$ for $K\geq 233$ and $p_{j}\equiv 1$
$\geq 2$, $<4$ $\geq 2$, $<K+1$
4+ $\geq 3$ $=2$ for $p_{j}\equiv 1$ $=m$ for $K$ sufficiently large
and $p_{j}\equiv 1$
Table 1: Overview of the competitive ratios obtained by our results. Recall that on $m$ machines, any algorithm is trivially $m$-competitive.
1.2 Preliminaries
Throughout this paper, for $n\in\mathbb{N}$, we denote by $[n]\coloneqq\{1,\ldots,n\}$ the set of numbers $1$ through $n$. An instance $I=(n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ of Online Makespan Scheduling under Scenarios$(m,K)$ (OMSS$(m,K)$) is parametrized by a number $m\in\mathbb{N}$ of machines and a number $K\in\mathbb{N}$ of scenarios, and defined as a tuple consisting of a number $n\in\mathbb{N}$ of jobs, a processing time $p_{j}\geq 0$ for each job $j$ and job subsets $S_{1},\ldots,S_{K}\subseteq[n]$, called scenarios. As we usually work with a single instance at a time, we often suppress this notation in the sequel as well as the parameters $m$ and $K$, which are often specified per section.
For a set $S\subseteq[n]$, we denote $p(S)\coloneqq\sum_{j\in S}p_{j}$. Our objective is to find a partitioning $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ of the set of jobs that minimizes $\max_{k\in[K]}\max_{i\in[m]}p(J_{i}\cap S_{k}).$ In other words, we consider schedules that arise from restricting a partition to each of the scenarios, and our goal is then to minimize the makespan of the worst-case scenario, which we also call the makespan of the schedule. Furthermore, our instance is online in the sense that parameters $m$ (number of machines) and $K$ (the number of scenarios) are known in advance, but jobs $j\in[n]$ as well as their processing times $p_{j}$ and the indices $k\in[K]$ such that $j\in S_{k}$ are revealed one job $j$ at a time. Once the information of a job $j$ is revealed in addition to already known information about the jobs in $[j-1]$, the job $j$ must be assigned to a machine $i\in[m]$, which is an irrevocable decision.
In our analyses and arguments, we use the notion of a completion time of a job $j$. In classical scheduling, this is defined as the sum of processing times preceding (including) $j$ that are on the same machine as $j$. In our setting, the restricted schedules are evaluated without idle times, meaning that such completion times might be different for different scenarios. Accordingly, we define the completion time of job $j\in J_{i}\cap S_{k}$ in scenario $S_{k}$ as $C^{k}_{j}\coloneqq\sum_{j^{\prime}\in J_{i}\cap S_{k}\cap[j]}p_{j^{\prime}}$ and the completion time as $C_{j}\coloneqq\max_{k\in[K],j\in S_{k}}C_{j}^{k}$.
Online algorithms are often analyzed by their performance compared to the best possible outcome under full information. In existing literature, this comparison is quantified by the competitive ratio. An algorithm is said to have competitive ratio $\rho$ if for every instance, it is guaranteed to output a solution that has an objective value which deviates from the value of an optimal solution under full information by at most a factor of $\rho$.
We also consider the special case of unit processing times $p_{j}\equiv 1$. In this special case, we may view our jobs as nodes, our scenarios as hyperedges containing a subset of the nodes and the partition $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ as a coloring $c\colon[n]\to[m]$ of the nodes. Then, our problem can be formulated as an online hypergraph coloring problem, in which we color a hypergraph $H=(V,E)$ with a fixed number of colors $1,\ldots,m$ and seek to minimize $\max_{i\in[m],e\in E}|\{c^{-1}(i)\}\cap e|$. The nodes $j\in[n]$ are revealed one at a time in increasing order along with partial hyperedges $e\cap[j]$ $(e\in E,j\in[n])$. [^2] According to this definition, hyperedges can have size $1$, although such hyperedges can and will be omitted as they pose no difference with respect to the objective function. We shall name this problem Online Makespan Hypergraph Coloring$(m)$ (OMHC$(m)$) and address it in Section 4.
1.3 Related Work
The problem that we introduce in this paper lies in the intersection of several well-known problems, which we would like to mention briefly.
Optimization under Scenarios.
In recent years, scenario-based models have received significant attention. An instance of a scenario model over a discrete optimization problem is often given as an instance of the underlying problem; and in addition, subsets of the underlying instance (called scenarios) are specified. The objective function involves the objective values that are attained restricted to each scenario; e.g., the maximum of these values over scenarios or their average.
Although a significant portion of the literature is rather recent, some well-known problems have been scenario-based problems in disguise. For instance, the famous exact matching problem ([exactmatching]), which asks whether a given graph whose edges are colored blue or red has a perfect matching with exactly $k$ blue edges for a given $k$, is equivalent to the matching problem under two scenarios, whereas a stochastic scenario-based model for matchings has recently been studied in [stocscenmatching]. Other examples include bin packing ([binpacking][binpacking2]), metric spanning tree ([mst]) and the traveling salesperson problem ([Ee2018priori]).
Among the scenario-based problems with the widest array of studies is machine scheduling with its many variations, see [shabtay2022] for an extensive overview. The closest to our problem is [feuerstein2016minimizing], in which the worst-case makespan is being minimized for parallel machine scheduling on two machines and a variable number of scenarios. Although an instance under a variable number of scenarios is not approximable with a ratio better than the trivial bound of $2$, a constant number of scenarios allows a PTAS.
In some variants of scenario scheduling, e.g., in [bosman2023total], a complexity gap is observed between two and three scenarios. Interestingly, we obtain a distinction between two and three scenarios as well, at least when we restrict to the special case of two machines.
Online Scheduling.
As it is impossible to capture the vast amount of literature on all online scheduling problems, let us restrict to online parallel machine scheduling with the objective of minimizing the makespan. This is precisely the special case $K=1$ of our problem.
The problem was first addressed in Graham’s seminal work [graham] from which we also draw inspiration in multiple ways. The competitive ratio of $2-\nicefrac{{1}}{{m}}$ for $m$ machines was not beaten for three decades, until it was improved slightly in [galamboswoeginger]. Graham’s List Scheduling Algorithm is evidently best possible for $m=2$, and it was observed to be best possible for $m=3$ as well [grahamlb]. In the following years, a stream of research narrowed the gap between upper and lower bounds for competitive ratios depending on $m$ ([sched1][sched2][sched3]), although for $m\geq 4$, still no tight bound is known. The state-of-the-art results are Fleischer and Wahl’s $1.9201$-competitive algorithm ([fleischerwahl]) and an asymptotic lower bound of $1.88$ due to Rudin III ([rudinphd]). For $m=4$, the best known lower bound is $\sqrt{3}\approx 1.732$, due to Rudin III and Chandrasekaran ([m4sqrt3]).
Discrepancy Minimization, Hypergraph Coloring.
A problem similar to Online Makespan Hypergraph Coloring, often called Online Hypergraph Coloring, has been studied extensively over the years. In this problem, the objective is to use the smallest number of colors while avoiding a monochromatic hyperedge, i.e., a hyperedge whose nodes all attain the same color. For hypertrees with $n$ nodes, an $O(\log(n))$-competitive algorithm is known to be best possible ([hypertreecoloring]), while for general graphs with $n$ nodes, no $o(n)$-competitive algorithm can exist ([NAGYGYORGY200823]). However, this standard model differs from ours, as a hyperedge is revealed only once all of its nodes have been revealed. This nuance deems it much easier to construct families of instances to achieve lower bounds. Moreover, due to differences in their respective objective functions, this problem does not seem to have a direct relation to Online Makespan Hypergraph Coloring. However, our approach has implications for a variant of Online Hypergraph Coloring in which partial hyperedges are known (see Section 1.1).
The reader may also notice similarities to the well-known discrepancy minimization problem (see e.g. [sixstddeviations][tightdisc]): Indeed, in the usual setting of discrepancy minimization, we have $m=2$ and the objective is to minimize the difference $\max_{e\in E}||\{c^{-1}(1)\}\cap e|-|\{c^{-1}(2)\}\cap e||.$ If all hyperedges $e\in E$ have the same size, this objective function is even equivalent to ours. In this regard, the problem that we study generalizes the $\ell_{\infty}$ norm online discrepancy minimization [onlinedisc] in uniform set systems as well.
2 Extending Graham’s List Scheduling
Our first observation is a simple algorithm that generalizes Graham’s List Scheduling Algorithm for OMSS$(m,K)$ and performs well when there are too many machines to force the schedule to create significant discrepancy. Our algorithm beats the trivial bound of $m$ if $m>K+1$.
We first observe that if the number $m$ of machines is large enough, we can apply the pigeonhole principle to find a machine that is not among the most loaded machines of any scenario. For $s\in[m-1]$ and $k\in[K]$, we call a machine $i$ $s$-favorable with respect to the $k$-th scenario if there are at least $s$ distinct machines $i^{\prime}\neq i$ that each have at least the load of machine $i$ in the $k$-th scenario, i.e., $p(J_{i^{\prime}}\cap S_{k})\geq p(J_{i}\cap S_{k})$.
Observation 2.1.
There exists a machine $i$ that is $\left(\left\lceil\frac{m}{K}\right\rceil-1\right)$-favorable with respect to every scenario $k\in[K]$.
Proof 2.2.
There are at most $s$ machines per scenario that are not $s$-favorable. Let $s=\left\lceil\frac{m}{K}\right\rceil-1$. Since $m>s\cdot K$, there must be a machine $i$ which is not non-$s$-favorable with respect to any scenario.
In light of this observation, we propose the following algorithm: Upon the assignment of each job $j$, we find a machine $i$ that is $\left(\left\lceil\nicefrac{{m}}{{K}}\right\rceil-1\right)$-favorable with respect to all scenarios $k$ with $j\in S_{k}$ and assign the job $j$ to the machine $i$.
Theorem 2.3.
The above algorithm is $\left(\frac{m-1}{\left\lceil\frac{m}{K}\right\rceil}+1\right)$-competitive for $m>K$.
Notice that the aforementioned competitive ratio is strictly smaller than $K+1$.
This algorithm is also applicable in the special case where every job appears in at most $k$ scenarios for a fixed $k\in\mathbb{N}$, yielding a ratio of $\nicefrac{{(m-1)}}{{\left\lceil\frac{m}{k}\right\rceil}}+1$. Furthermore, plugging in $K=1$ yields the well-known competitive ratio of $2-\nicefrac{{1}}{{m}}$ of Graham’s List Scheduling Algorithm.
3 Near-Tight Bounds for $m=2$ Machines
When the partitioning is among two machines, a clear line can be drawn: Any algorithm is trivially $2$-competitive, and if we have at least $3$ scenarios, this is best possible.
Observation 3.1.
For $m=2$ and $K\geq 3$, there exists no $r$-competitive algorithm for any $r<2$, even for the special case of unit processing times.
Proof 3.2.
For all jobs $j$ below, we assume $p_{j}=1$. The first instance $\mathcal{I}_{1}$ is given as follows: $n=m+1$, $S_{1}=\{1,3,\ldots,m+1\}$, $S_{2}=\{2,3,\ldots,m+1\}$.
We consider an arbitrary algorithm that places the first jobs $1$ and $2$ to two distinct machines, without loss of generality, to the first and second, respectively. In this case, the makespan at the end must be $2$, as two jobs $(j,j^{\prime})\neq(1,2)$ must be assigned to the same machine. However, the following assignment, which places the first two jobs together would have yielded a completion time of $1$: $J_{1}=\{1,2\},J_{i}=\{i+1\}\text{ for }2\leq i\leq m.$ Hence, no algorithm that places two unit weight jobs in $S_{1}\setminus S_{2}$ resp. $S_{2}\setminus S_{1}$ to two separate machines can attain a competitive ratio better than $2$.
In light of this, we consider another instance $\mathcal{I}_{2}$, on which assigning the first two jobs to the same machine leads to a failure. It is given by $n=m+2$, $S_{1}=\{1,3\}$, $S_{2}=\{2,4,5,6,\ldots,m+2\}$, $S_{3}=\{3,4,5,\ldots,m+2\}$. The first two jobs are indeed identical to those of $\mathcal{I}_{1}$.
By our considerations about the instance $\mathcal{I}_{1}$, we may assume that the first two jobs are assigned to the same machine, the first one without loss of generality. However, any assignment that places the first two jobs to the same machine admits makespan $2$: Otherwise, due to the third scenario, the jobs $\{3,\ldots,m+2\}$ must occupy one machine each. The job $j$ that is placed to the first machine has a common scenario with either job $1$ (if $j=3$), or with job $2$ (otherwise).
On the other hand, this statement cannot be extended to $K=2$. To show this, we present a minimalistic yet insightful algorithm (Algorithm 1) for makespan scheduling under $K=2$ scenarios and $m=2$ machines. Our general idea is to fix a simple rule for assigning double-scenario jobs, i.e., jobs in $S_{1}\cap S_{2}$, and partition the single-scenario jobs $j\in S_{1}\triangle S_{2}$ so that the machines are reasonably balanced even after upcoming hypothetical double-scenario jobs.
Rule 3.3.
If $j\in S_{1}\cap S_{2}$, then assign the job $j$ to a machine $i$ such that the makespan of the schedule was previously attained by the other machine $3-i$. In case of a tie, select the machine that received the job $j-1$ ($j=1$ is assigned to the first machine if applicable).
Throughout our analysis, we only use two types of lower bounds for the offline optimum: Processing times $p_{j}$ of selected jobs $j$ whose completion times equal the makespan, and the average load of a selected scenario $S_{k}$. Hence, while searching for the worst-case sequence of the upcoming double-scenario jobs, we may replace the actual definition of competitiveness with that of a proxy one:
Definition 2.
We define the proxy competitive ratio of a schedule as
$\rho=\frac{\max_{i,k\in\{1,2\}}p(S_{k}\cap J_{i})}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k})}{2}\right\},\max_{j\in[n]\cap\mathrm{argmax_{j\in[n]}\{C_{j}\}}}\{p_{j}\}\right\}},$
where $C_{j}$ denotes the completion time of the job $j$.
The proxy competitive ratio is an upper bound on the competitive ratio. The lower bounds that we use are already well-known; e.g., Graham’s List Scheduling Algorithm is analyzed by bounding the proxy competitive ratio from above. We take this approach, however, one step further and restrict ourselves to analyzing only some instances that attain the largest possible proxy competitive ratio (see next Observation). This is significantly more difficult for the competitive ratio in the usual sense, as a witness to having good or bad offline optima highly depends on how the processing times of jobs fit together numerically.
Observation 3.4 (cf. Figure 1).
Let $\mathcal{A}$ be an online algorithm for OMSS(2,2)22(2,2)( 2 , 2 ) obeying Rule 3.3. Assume that $\mathcal{A}$ has a proxy competitive ratio of at most $\frac{5}{3}$ on instances where
  1. (i)
    $n\in S_{1}\cap S_{2}$, and
  2. (ii)
    $\mathcal{A}$ outputs a schedule with the following property: If $j\in[n]$ is the largest index with $j\in S_{1}\triangle S_{2}$, then $C_{j}=C_{n-1}$.
Then, $\mathcal{A}$ is $\frac{5}{3}$-competitive.
The above observation motivates the following parameter of significant value.
[…]
Figure 1: Illustrative description of Observation 3.4. Processing times $p_{j}$ are given by the widths. Given the dark blue schedule, a worst outcome turns out to be the light blue job $n-1$ with a processing time $p_{n-1}$ such that $C_{n-1}=C_{j}$, followed by the light red job $n$ with $p_{n}=p(S_{1}\cap[n-1])$.
Definition 3.
Consider a schedule where, without loss of generality, the makespan is attained at $S_{1}\cap J_{1}$. The anticipation $\alpha$ of this schedule is given by the ratio
$\alpha\coloneqq\begin{cases}\frac{p(J_{1}\cap S_{1})}{\max\left\{p(J_{1}\cap S_{2}),p(J_{2}\cap S_{1})+p(J_{1}\cap S_{1})-p(J_{2}\cap S_{2})\right\}}&\text{if }p(J_{2}\cap S_{2})>p(J_{2}\cap S_{1}),\\ 0&\text{ else.}\end{cases}$
By convention, we assume $\frac{0}{0}=1$.
In Figure 1, we have $p(J_{2}\cap S_{2}\cap[n-2])>p(J_{2}\cap S_{1}\cap[n-2])$ and the second term of the denominator corresponds to the load $p(J_{2}\cap S_{1}\cap[n-1])$, i.e., the load of $J_{2}\cap S_{1}$ after the $(n-1)$-st job is assigned. In general, anticipation measures the minimum discrepancy of a machine with respect to different scenarios after machine $2$ is loaded with double-scenario jobs just enough to attain the makespan of the schedule together with machine $1$. In our algorithm, one of the goals is to keep $\alpha\leq 2$ in order to brace ourselves for upcoming double-scenario jobs. The invariant which we maintain to this end is the following.
Invariant 3.5.
The schedule satisfies the following assertions:
  1. (i)
    Its proxy competitive ratio is bounded by $\rho\leq\nicefrac{{5}}{{3}}$.
  2. (ii)
    The schedule has anticipation bounded by $\alpha\leq 2$.
  3. (iii)
    Let $k,i$ be given so that $J_{i}\cap S_{k}$ admits the makespan in the schedule. If the schedule is dominated by the $i$-th machine, i.e., $p(J_{i}\cap S_{k^{\prime}})>p(J_{3-i}\cap S_{k^{\prime\prime}})$ for all $k^{\prime},k^{\prime\prime}\in\{1,2\}$, then we have $p(J_{i}\cap S_{k})\leq 2p(J_{i}\cap S_{3-k})$.
The last property might appear somewhat unnatural at the first glance, though it is crucial for bounding the anticipation $\alpha\leq 2$ in the subsequent iterations.
Algorithm 1 shows how a single job $j$ is assigned given a schedule of the first $j-1$ jobs where Invariant 3.5 is maintained. To analyze the algorithm, we prove that instances satisfying the properties in Observation 3.4 are always assigned so that the first $n-1$ jobs maintain Invariant 3.5 and the last job results in a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$.
Algorithm 1 The assignment of a single job $j$ given a $\frac{5}{3}$-competitive schedule that satisfies Invariant 3.5.
1:Rename machines and scenarios so that $(J_{1}\cap S_{1})$ is maximal
2:if $j\in S_{1}\setminus S_{2}$ then
3:     return 2
4:end if
5:if $j\in S_{2}\setminus S_{1}$ then
6:     if assigning $j$ to the first machine satisfies Invariant 3.5 then
7:         return 1
8:     else
9:         return 2
10:     end if
11:end if
12:if $j\in S_{1}\cap S_{2}$ then
13:     return $\mathrm{argmin}_{i=1,2}\max_{k=1,2}\{p(J_{i}\cap S_{k})\}$
14:end if
A significant portion of the analysis is focused on proving that if line 9 is executed, Invariant 3.5 is indeed maintained. We refer the reader to the full version of the paper for further details.
In spite of the restrictions that arise from both the fixed double-scenario rule and how we evaluate lower bounds, we find out that our algorithm in the previous subsection is not far from being best possible.
Theorem 3.6.
For $m=2$ and $K=2$, there exists no deterministic algorithm with competitive ratio strictly smaller than $\frac{9+\sqrt{17}}{8}\approx 1.640$.
Proof 3.7.
To ease the notation once again, let $X\coloneqq 3+\sqrt{17}$.
The first four jobs are given by $1,2\in S_{1}\setminus S_{2}$, $3,4\in S_{2}\setminus S_{1}$, $p_{1}=p_{2}=p_{3}=p_{4}$. Any algorithm with competitive ratio better than $2$ must assign these jobs so that the makespan equals $1$ at the end; otherwise, we may stop revealing further jobs so that the offline optimum equals $1$, while the makespan equals $2$.
The fifth job fulfills $5\in S_{1}\setminus S_{2}$ and $p_{5}=X$. Without loss of generality, it is assigned to the first machine.
The sixth job is given by $6\in S_{2}\setminus S_{1}$, $p_{6}=\frac{X}{2}$. Now there are two options:
  1. (i)
    If the sixth job is also assigned to the first machine, we reveal the seventh job with $7\in S_{2}\setminus S_{1}$ and $p_{7}=X$. This job cannot be assigned to the first machine, as the makespan would then be $1+\frac{3X}{2}$, while the offline optimum is $X$. Indeed, this would imply a ratio of at least
    $\frac{1+\frac{9+3\sqrt{17}}{2}}{3+\sqrt{17}}=\frac{9+\sqrt{17}}{8}.$
    In particular, we obtain $p(S_{1}\cap J_{1}\cap[7])=p(S_{2}\cap J_{2}\cap[7])=1+X$.
  2. (ii)
    If the sixth job is assigned to the second machine, we reveal the seventh job as $7\in S_{1}\cap S_{2}$ and $p_{7}=\frac{X}{2}$. Due to similar reasoning to the other case, the seventh job must be assigned to the second machine and we again have $p(S_{1}\cap J_{1}\cap[7])=p(S_{2}\cap J_{2}\cap[7])=1+X$.
See Figure 2 for the first seven jobs in each of the cases.
In both of these cases, we reveal the eighth job with $8\in S_{1}\cap S_{2}$ and $p_{8}=2+\frac{3X}{2}$. No matter where this job is assigned, the makespan is $3+\frac{5X}{2}$, whereas the optimal solution $\{1,\ldots,7\}\dot{\cup}\{8\}$ achieves a makespan of $2+\frac{3X}{2}$. Therefore, the competitive ratio is bound to be at least
$\frac{3+\frac{5X}{2}}{2+\frac{3X}{2}}=\frac{6+5(3+\sqrt{17})}{4+3(3+\sqrt{17})}=\frac{9+\sqrt{17}}{8},$
as desired.
[…]
Figure 2: The first seven jobs in cases (i) and (ii) of the proof of Theorem 3.6, respectively. In both cases, an eighth job $8\in S_{1}\cap S_{2}$ forces the schedules to the desired lower bound of competitiveness.
Notice that the aforementioned family of instances fails to achieve small competitive ratios for reasons similar to those that require Property (iii) in Invariant 3.5.
It can also be shown through a simple array of instances that, no algorithm that satisfies Rule 3.3 can beat the bound of $\nicefrac{{5}}{{3}}$, for which we refer the reader to the full version of the paper.
4 Tight Bounds for Sufficiently Many Scenarios
One of our main results (Theorem 1.2) states that for any number $m\in\mathbb{N}$ of machines, there exists a number $K$ of scenarios such that no deterministic algorithm for OMSS$(m,K)$ can have a competitive ratio less than the trivial upper bound $m$, even for unit processing times.
Due to space constraints, we merely sketch a proof for the special case of $m=3$. While doing so, we showcase key ideas of the general argument, although we forfeit the restriction onto hypertrees for the sake of a more clearly structured description. In the context of scheduling, this is indeed not a significant change.
Theorem 4.1.
For OMHC$(3)$, there exists no $r$-competitive online algorithm for any $r<3$, even when restricted to instances with at most $233$ hyperedges of size at most $3$ each.
The problem OMHC$(m)$, which we formulated in Section 1.2, is equivalent to the unit processing time case of OMSS$(m,K)$ with $K$ part of the input, so that the lower bounds are immediately implied for the latter as well.
To prove Theorem 4.1, we provide a family of instances on which any online algorithm is forced to output a monochromatic hyperedge of size $3$. For each of the instances, we also maintain (offline) colorings $c\colon[j]\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$, with the property that if two nodes $v,v^{\prime}$ are contained in a common hyperedge, we have $c(v)\neq c(v^{\prime})$. The latter coloring suggests an offline optimal value of $1$ for each of the instances, implying the competitive ratio of $3$ immediately. The main difference between the online and offline colorings is, as expected, that we may change the offline coloring completely at all times as long as its makespan equals $1$. Among other modifications, we often employ a permutation $\pi\colon\{\mathrm{red,blue,yellow}\}\to\{\mathrm{red,blue,yellow}\}$ on connected components to realize this possibility. We must trace the offline coloring simultaneously because the hyperedges are often defined adversarially according to both colorings of the current hypergraph.
To differentiate between the online and offline colorings, we denote by $J_{i}\subseteq V$ the subset of nodes assigned to the online color $i$ for $i\in\{1,2,3\}$. Right before the revelation of a node $j$, we implicitly mean $J_{i}\cap[j-1]$ whenever we mention $J_{i}$ for some $i\in\{1,2,3\}$. Inspired by our original scheduling problem, we call the objective value $\max_{e\in E}\max_{i\in\{1,2,3\}}|e\cap J_{i}|$ of an instance its makespan. Without loss of generality, we truncate an instance as soon as the online solution is forced to obtain a makespan of $3$.
Our gadget: Seven-node subhypergraph $\mathcal{S}$.
Initially, we study a subhypergraph $\mathcal{S}$ that we use as a building block throughout our construction. The hypergraph $\mathcal{S}$ is given by the nodes $V(\mathcal{S})=\{1,\ldots,7\}$ and hyperedges
[…]
The hyperedges have size at most $3$ and the set $E(\mathcal{S})$ is closed under taking subsets of size at least $2$, i.e., if $e\in E(\mathcal{S})$, $|e|=3$ and $e^{\prime}\subseteq e$ with $|e^{\prime}|=2$, then $e^{\prime}\in E(\mathcal{S})$. The hypergraph $\mathcal{S}$ together with the maximal hyperedges $e_{1},\ldots,e_{5}$ is given in Figure 3.
$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$3333111122225555444466667777
Figure 3: The hypergraph $\mathcal{S}$ with the maximal hyperedges $e_{1},\ldots,e_{5}$.
We first assume that the seven nodes in $V(\mathcal{S})$ are assigned to exactly two out of the three online colors. Since the online colors are initially symmetric, let us assume that we use the online colors $1$ and $2$. By a careful case distinction, we may observe:
Proposition 4.
For any partitioning $V(\mathcal{S})=J_{1}\dot{\cup}J_{2}$ of the nodes that admits a makespan of at most $2$, there exists a node coloring $c\colon V(\mathcal{S})\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ such that there are two hyperedges $e\in E(\mathcal{S}[J_{1}]),e^{\prime}\in E(\mathcal{S}[J_{2}])$ with $c(v)\in\{\mathrm{red},\mathrm{blue}\}$ (or another subset of $\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ of size $2$) for all $v\in e\cup e^{\prime}$.
In other words, we have induced monochromatic hyperedges of size $2$ with two different colors, both with respect to the online coloring. Applying this repeatedly, we can further force the algorithm to a monochromatic hyperedge of size $3$, which we show next.
Pigeonholing the two-color case.
In our construction, we would like to create copies of the graph $\mathcal{S}$ successively, which we denote by $\mathcal{S}(t)$ for incrementing $t\in\mathbb{N}$.
A straightforward application of the pigeonhole principle yields that, if we have seven subhypergraphs $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ that are each partitioned onto two online colors, at least three of them (without loss of generality $\mathcal{S}(1)$, $\mathcal{S}(2)$, $\mathcal{S}(3)$) must be partitioned onto the same two online colors. Again without loss of generality, these two online colors are $1$ and $2$. Considering Proposition 4, we can permute the colorings of the three copies $\mathcal{S}(1)$, $\mathcal{S}(2)$ and $\mathcal{S}(3)$ each offline so that for any offline color $u\in\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$, there exists a copy $\mathcal{S}(t)$ with two (online) monochromatic hyperedges $e_{i}^{t}\in E(\mathcal{S}(t))$ $(i\in\{1,2\})$ of size $2$ avoiding the color $u$. The hypergraph in Figure 4 restricted to nodes of the form $j_{s}^{i}$ illustrate these edges.
Once the six edges in question are found, one can insert nodes $x_{1}$, $x_{2}$, $x_{3}$ in this order, revealing hyperedges as depicted in Figure 4. To avoid a monochromatic edge, both $x_{1}$ and $x_{2}$ must be assigned the third online color. Keeping this in mind, it follows that $x_{3}$ cannot be assigned an online color without creating a monochromatic edge with respect to the online coloring. Together with the fact that the offline coloring indeed has makespan $1$, we obtain:
Lemma 4.2.
Any algorithm that assigns at least $7$ copies of $\mathcal{S}$ to at most two online colors each is no better than $3$-competitive.
[…]
Figure 4: All fifteen relevant nodes together with the hyperedges within them. The heights of the vertices encode their online color ($J_{1}$, $J_{2}$, $J_{3}$), while their printed colors denote their offline color. Any assignment of the node $x_{3}$, given the other nodes and induced hyperedges, leads to a makespan of $3$.
Subhypergraph $\mathcal{S}$ on Three Online Colors: The Palettes.
It remains to consider the case where copies of the subhypergraph $\mathcal{S}$ are partitioned among the three colors, i.e., its nodes $\{1,\ldots,7\}$ are partitioned in nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$.
Lemma 4.3.
For every online coloring $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$ of the nodes $V(\mathcal{S})=\{v_{1},\ldots,v_{7}\}$ for nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$, there is an offline coloring $c\colon V(\mathcal{S})\to\{\mathrm{red},\mathrm{blue},\mathrm{yellow}\}$ such that two vertices exist with different online colors but the same offline color.
In our construction, we exploit this property as follows: Let us consider a copy $\mathcal{S}(t)$ and fix two of its nodes $v_{i}(t)\in J_{i}$, $v_{i^{\prime}}(t)\in J_{i^{\prime}}$ provided by Lemma 4.3 as well as a third node $v_{i^{\prime\prime}}(t)\in J_{i^{\prime\prime}}$, where $i^{\prime\prime}\notin\{i,i^{\prime}\}$. By their choice, the nodes $v_{i}(t)$ ($i\in\{1,2,3\}$) admit at most two offline colors and avoid a third offline color $C\in\{\mathrm{blue},\mathrm{red},\mathrm{yellow}\}$.
In this case, the subhypergraph $\mathcal{S}(t)$ is called a $C$ palette (i.e., yellow palette, blue palette or red palette) and the corresponding nodes $v_{i}(t)$ are called palette nodes. In particular, $c(v_{i}(t))\neq C$ holds for the palette nodes of a $C$ palette. The choice of the name is motivated by the fact that, if we connect a new node $w$ to e.g. yellow palette nodes, we are still allowed to extend the offline coloring by $c(w)=\mathrm{yellow}$, while maintaining that the makespan of the offline coloring is $1$.
The Construction.
Below is a summary of our construction for the case where Lemma 4.2 is not applicable. Essentially, we create copies of the subhypergraph $\mathcal{S}$ which we connect with the existing hypergraph. The online color of the node $v$ is denoted by $\varphi(v)$ and its (current) offline color by $c(v)$.
  1. (i)
    We create up to $13$ copies of the subhypergraph $\mathcal{S}$. By Lemma 4.2, we may assume that at least $7$ of the copies are assigned to all three online colors (we truncate this step as soon as $7$ such copies can be found). We handpick $7$ such copies of $\mathcal{S}$, calling them $\mathcal{S}(1),\ldots,\mathcal{S}(7)$.
  2. (ii)
    We permute the offline coloring on each copy of $\mathcal{S}$ so that $\mathcal{S}(1)$, $\mathcal{S}(4)$ and $\mathcal{S}(7)$ are yellow palettes, $\mathcal{S}(2)$ and $\mathcal{S}(5)$ are blue palettes, and $\mathcal{S}(3)$ and $\mathcal{S}(6)$ are red palettes.
  3. (iii)
    We reveal a node $w_{\mathcal{A}}$ which is connected to palette nodes $v_{i}(1)\in V(\mathcal{S}(1))\cap J_{i}$ for $i\in\{1,2,3\}$. The assignment to the online color $\varphi(w_{\mathcal{A}})$ leads to a monochromatic hyperedge $\{v_{\varphi(w_{\mathcal{A}})}(1),w_{\mathcal{A}}\}$ of online color $\varphi(w_{\mathcal{A}})$. We define $v_{\mathcal{A}}\coloneqq v_{\varphi(w_{\mathcal{A}})}(1)$.
  4. (iv)
    We reveal three nodes $w_{\mathcal{B}1}$, $w_{\mathcal{B}2}$ and $w_{\mathcal{B}3}$ that are connected by the hyperedge $\{w_{\mathcal{B}1},w_{\mathcal{B}2},w_{\mathcal{B}3}\}$ to each other. Moreover, for each $s\in\{1,2,3\}$, the node $w_{\mathcal{B}s}$ is connected to the palette nodes $v_{i}(s+1)$ for $i\in\{1,2,3\}$. There is an $i\neq\varphi(w_{\mathcal{A}})$ such that $\{v_{i}(s+1),w_{\mathcal{B}s}\}$ is a monochromatic edge of online color $i$ for a suitable copy $\mathcal{S}(s)$, $s\in\{1,2,3\}$. For this $s$, we define $v_{\mathcal{B}}\coloneqq v_{i}(s+1)$ and $w_{\mathcal{B}}\coloneqq w_{\mathcal{B}s}$.
[…]
    Figure 5: An example constellation of the three palettes together with the nodes $w_{\mathcal{B}t}$. The $y$-axis in the palettes denote the online colorings of the palette nodes respectively. The offline colors of the palette nodes might look differently, but they are never the same color as the palette, which is depicted as the color of the rectangles. The online coloring of the nodes $w_{\mathcal{B}i}$ are not yet known.
  5. (v)
    We reveal further copies of $\mathcal{S}$ together with edges described below, until a copy $\mathcal{C}$ is assigned to all three online colors. In the copy $\mathcal{C}$, we denote the nodes $v_{i}(t)$ by $w_{\mathcal{C}t}$ for clarity. The nodes $w_{\mathcal{C}t}$ of the copy $\mathcal{C}$ are connected to the palettes $\mathcal{S}(5)$, $\mathcal{S}(6)$, $\mathcal{S}(7)$ by edges depending on their offline colors, in that nodes $w_{\mathcal{C}t}$ with $c(w_{\mathcal{C}t})=C$ are connected to the palette nodes $v_{i}(s)$ for all $i\in\{1,2,3\}$ and $s\in\{5,6,7\}$ such that $\mathcal{S}(s)$ is a $C$ palette, see Figure 6. As the nodes $w_{\mathcal{C}t}$ are assigned to all three online colors, a makespan of $2$ is also achieved in the online color $i\notin\{\varphi(w_{\mathcal{A}}),\varphi(w_{\mathcal{B}})\}$. More precisely, there exists a monochromatic edge $\{v_{i}(s),w_{\mathcal{C}t}\}$ of the online color $i$. We rename $v_{\mathcal{C}}\coloneqq v_{i}(s)$ and $w_{\mathcal{C}}\coloneqq w_{\mathcal{C}t}$.
  6. (vi)
    We permute the offline coloring in each of the connected components created in (iii), (iv) and (v) again, so that $c(v_{j})=\mathrm{blue}$ and $c(w_{j})=\mathrm{yellow}$ for $j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$.
  7. (vii)
    We add a final node $v_{n}$ together with hyperedges $\{v_{n},v_{j},w_{j}\}$ for each $j\in\{\mathcal{A},\mathcal{B},\mathcal{C}\}$. Now, the online color $\varphi(v_{n})$ that this node is assigned to admits a makespan of $3$, while extending the offline coloring by $c(v_{n})=\mathrm{red}$ maintains an offline coloring with makespan $1$.
[…]
Figure 6: Partial sketch of subinstance $\mathcal{C}$.
Notice that we have maintained an offline coloring of the nodes with makespan $1$ so far up to the very last node, and the last node $n$ neighbors only blue and yellow nodes by construction and therefore the offline assignment $c(n)=\mathrm{red}$ does not increase the makespan. In total, this yields that the offline optimal value is $1$. A careful enumeration of introduced hyperedges reveals that $233$ hyperedges suffice to create a monochromatic hyperedge for any of the cases demonstrated, which proves Theorem 4.1.
A Note on Theorem 1.2.
Here, we abstain from presenting a proof of Theorem 1.2. Our construction in this general case is also based on incrementally creating monochromatic edges of each online color. However, we are faced with two fundamental challenges:
  1. (i)
    The monochromatic hyperedges increase in size rather slowly, yet unless they have size of $m-1$, they are not strong enough to act as a palette for other hyperedges in the same sense as here.
  2. (ii)
    To use our construction for Corollary 1, we restrict to hypertrees. This is a significant constraint, as it roughly implies that every palette may be used at most once.
Therefore, we include a more intricate recursive construction with much larger hypergraphs in the full version of the paper.
5 Tight Bounds For $K=3$ and Unit Processing Times
In the previous section, we have established that for large enough $K$, there is no better-than-$3$-competitive algorithm for OMSS$(3,K)$, even for unit processing times. Let us reverse the parameters and ask: How competitively can we solve OMSS$(m,3)$? This is not an easy question, indeed, even for OMSS$(m,1)$, $m\geq 4$, a tight bound is not known. Restricting to the unit processing time case $(p_{j}\equiv 1)$, we again obtain a tight bound, which is the smallest possible ratio (due to Theorem 2.3):
Theorem 5.1.
For all $m\in\mathbb{N}$, there is a $2$-competitive algorithm for the special case of OMSS$(m,3)$ where the jobs have unit processing times.
The algorithm in question employs an adapted round-robin method which assigns jobs to machines according to the scenarios they appear in. Essentially, we try to assign jobs $j\in S_{1}\setminus(S_{2}\cup S_{3})$ and $j^{\prime}\in(S_{2}\cup S_{3})\setminus S_{1}$ in a similar manner, and analogously for cyclic permutations of the indices $i\in\{1,2,3\}$. This gives a partition of jobs $j\in[n]\setminus(S_{1}\cap S_{2}\cap S_{3})$ in three categories. The round robin assignments for each category are translated by roughly $\nicefrac{{m}}{{3}}$, and jobs $j\in S_{1}\cap S_{2}\cap S_{3}$ are handled separately.
The algorithm is best visualized by drawing an analogue to a bingo card, for which we refer the reader to the full version.
6 Conclusion and Outlook
Throughout, we have established several competitiveness results for Online Makespan Scheduling under Scenarios. The first main takeaway from our work is a competitiveness gap between two and at least three scenarios when we consider $m=2$ machines. This result draws a parallel to the known tractability results of several other problems, in which the line between easy and hard was drawn between two and three scenarios. Surely, the fact that the subsets of $\{1,2,3\}$ are not laminar, has been a deciding factor in our simple non-competitiveness result, as has been in several of the well-known results. Interestingly, restricting to the proxy competitive ratio as well as fixing the method of assigning double-scenario jobs essentially reduced the existence problem to the description of a polyhedron. We believe that similar techniques might prove useful for other cases and related problems as well.
We have further compared the behavior of competitiveness for increasing number $K$ of scenarios and increasing number $m$ of machines. In the special setting of unit processing times and $m=3$ machines, the contrast is evident: On one hand, there is a $2$-competitive algorithm for three scenarios and $m$ machines for all $m\in\mathbb{N}$, which matches the lower bound obtained readily for $m=2$. On the other hand, for $K$ sufficiently large, the trivial upper bound of $3$ cannot be beaten. That being said, $K+1$ is a strict upper bound as well for a problem on $K$ scenarios. To summarize, for large values of $m$ and small values of $K$, we would expect the possibilities for competitive ratios to be dominated by $K$.
In light of our results, we may raise several open questions: In addition to the obvious task of finding tight bounds for all possibilities $(m,K)$ of the numbers of machines and scenarios, exploring the competitiveness as $m$ approaches to infinity in the weighted processing time case remains interesting. However, these problems have not been resolved even for $K=1$, whence it is fair to assume challenges in further progress. Another interesting research direction is to introduce migration, as seen in e.g. [migration]. In this model, we are allowed to revoke a bounded size of decisions in hindsight. Similarly, instead of minimizing the worst-case scenario, one could look into minimizing the average scenario or maximum regret.

### v2

@@
 We prove tight and near-tight bounds, several of which are achieved through novel constructions.
-In particular, we leverage the interplay between the unit processing time case of our problem and the hypergraph coloring problem both ways: We use hypergraph coloring techniques to steer an adversarial family of instances proving lower bounds for our problem, which in turn leads to lower bounds for several variants of online hypergraph coloring.
+In particular, we leverage the interplay between the unit processing time case of our problem and the hypergraph coloring problem both ways: We use hypergraph coloring techniques to steer an adversarial family of instances proving lower bounds, which in turn leads to lower bounds for several variants of online hypergraph coloring.
 keywords:
@@
 Moreover, even special cases of online makespan scheduling under scenarios correspond to interesting questions in their own regard.
-The special case $p_{j}\equiv 1$ of unit processing times is related to some cases of discrepancy minimization as well as hypergraph coloring.
+The special case $p_{j}\equiv 1$ of unit processing times can be related to some cases of discrepancy minimization as well as hypergraph coloring.
 In the large array of previous research on combinatorial optimization under scenarios, it was shown for several offline problems that optimizing under a number of scenarios is computationally harder than their underlying single-scenario counterparts.
-Accordingly, our goal in this work is to gain a better understanding of the intersection of scenario-based models and competitiveness, in particular how competitiveness diminishes under a growing number of scenarios.
+Accordingly, our goal in this work is to gain a better understanding of the intersection of scenario-based models and competitiveness, in particular how competitiveness becomes harder under a growing number of scenarios.
 1.1 Our Contribution
@@
 Several of the entries are obtained via a generalization of Graham’s List Scheduling Algorithm which we extend using the pigeonhole principle.
-By a simple example, one can show that for any number $m\geq 2$ of machines and (at least) $3$ scenarios, no algorithm with a competitive ratio better than $2$ can exist, even for jobs with unit processing times.
+By a simple example, we show that for any number $m\geq 2$ of machines and (at least) $3$ scenarios, no algorithm with a competitive ratio better than $2$ can exist, even for jobs with unit processing times.
 To underline the competitiveness gap between two and three scenarios, we contrast the aforementioned result with an algorithm for two machines and two scenarios, which is one of our main results.
@@
 This algorithm as well as its analysis combine techniques well-known from algorithms for online makespan scheduling with novel ideas that concern the scenario-based properties of a schedule.
-We first alter the definition of competitive ratio by allowing only specific lower bounds for an offlie optimum, which allows us to make structural modifications and restrictions on worst-case instances.
 What sets our analysis apart from previous work is that we alternate between forward and reverse engineering: Assuming that our algorithm obeys some fixed rules, we explore what we may assume about a well-structured worst-case instance.
@@
 The remainder of the paper demonstrates a disparity between increasing the number $m$ of machines and increasing the number $K$ of scenarios.
-For a fixed but arbitrary $m$, we show that there exists a sufficiently large number $K(m)$ of scenarios and a family of instances with $K(m)$ scenarios, on which no deterministic online algorithm can achieve a ratio better than $m$.
+For a fixed but arbitrary $m$, we show that there exists a sufficiently large number $K(m)$ of scenarios and a family of instances with $K(m)$ scenarios on which no deterministic online algorithm can achieve a ratio better than $m$.
 Furthermore, we can even choose all instances to have unit processing times $p_{j}\equiv 1$.
 This result is particularly noteworthy, considering that any online algorithm on $m$ machines is trivially $m$-competitive: Indeed, the online output is at most the total load and an offline optimum is at least the average load, and these are by at most a factor of $m$ apart.
-To find the suitable number $K(m)$, we interpret the instances as hypergraphs, passing to a problem we call Online Makespan Hypergraph Coloring$(m)$ instead (cf.
+To find the suitable number $K(m)$, we interpret the instances as hypergraphs, passing to a problem we call Online Makespan Hypergraph Coloring(m)(m) instead (cf.
 Section 1.2).
@@
 Let $m\in\mathbb{N}$.
-There exists a number $K(m)$ such that for any $r<m$, no deterministic algorithm for Online Makespan Hypergraph Coloring$(m)$ is $r$-competitive, even when restricted to hypertrees with $K(m)$ hyperedges.
-The statement of Theorem 1.2 is best possible in the sense that the numbers $K(m)$ cannot be replaced by a global $K$ for which the incompetitiveness holds for all $m$ (cf. Theorem 2.3).
+There exists a number $K(m)$ such that for any $r<m$, no deterministic algorithm for Online Makespan Hypergraph Coloring(m)(m) is $r$-competitive, even when restricted to hyperforests with $K(m)$ hyperedges.
+The statement of Theorem 1.2 is best possible in the sense that the numbers $K(m)$ cannot be replaced by a global $K$ for which the incompetitiveness holds for all $m$ (cf. Corollary 2.5).
 The proof of Theorem 1.2 can be found in the full version of the paper.
@@
 We utilize a subhypergraph on $7$ nodes as a gadget and obtain a relatively compact construction, both in terms of the hypergraph and the case distinction that we apply to analyze it.
-[…]
+An advantage of our approach while proving Theorem 1.2 is that it only features hyperforests.
+More precisely, we implicitly show that for every $m\in\mathbb{N}$, there is a number $n\leq(2m+1)\uparrow\uparrow 5$ of nodes such that no deterministic online algorithm on $n$ nodes that uses $m$ colors can avoid a monochromatic hyperedge.
+In other words, in order to color a hyperforest with an arbitrary number $n$ of nodes online without monochromatic edges, we need $\Omega(\mathrm{slog_{5}}n)$ colors (cf.
+Preliminaries for notation), even if the restrictions $e\cap\{1,\ldots,j\}$ of hyperedges $e$ to revealed nodes are known upon the revelation of nodes $\{1,\ldots,j\}$.
+Here, the emphasis is on the additional knowledge of partial hyperedges: This makes an adversarial revelation of hyperedges more challenging and to the best of our knowledge, none of the previous lower bound constructions [NAGYGYORGY200823][hypertreecoloring] are applicable under such additional information.
+This result addresses a completely open problem from [NAGYGYORGY200823] and sets a non-constant lower bound for Online Hypergraph Coloring for hypertrees under the knowledge of partial hyperedges.
 Corollary 1.
@@
 m $\downarrow$ K $\rightarrow$ 2 3 4+
-2 $>1.640$, $\leq\nicefrac{{5}}{{3}}$ $=2$ $=2$
-3 $\leq 2$ $\geq 2$, $\geq 2$
-$=2$ for $p_{j}\equiv 1$ $=3$ for $K\geq 233$ and $p_{j}\equiv 1$
-$\geq 2$, $<4$ $\geq 2$, $<K+1$
-4+ $\geq 3$ $=2$ for $p_{j}\equiv 1$ $=m$ for $K$ sufficiently large
-and $p_{j}\equiv 1$
+2 $\in\left(1.640,{{\frac{5}{3}}}\right]$ =2{=2} =2{=2}
+3 $\in\left[\frac{5}{3},2\right]$}
+$[2,3]$
+=2{=2} for $p_{j}\equiv 1$}
+$\in[2,3]$ =3=3 even for $p_{j}\equiv 1$, if $K\geq 233$
+4+ $\in\left[\frac{5}{3},3\right)$
+$\in[2,4)$
+=2{=2} for $p_{j}\equiv 1$}
+$\in[2,\min\{m,K+1\})$ =m=m even for $p_{j}\equiv 1$, if $K$ large enough
 Table 1: Overview of the competitive ratios obtained by our results. Recall that on $m$ machines, any algorithm is trivially $m$-competitive.
 1.2 Preliminaries
-Throughout this paper, for $n\in\mathbb{N}$, we denote by $[n]\coloneqq\{1,\ldots,n\}$ the set of numbers $1$ through $n$.
-An instance $I=(n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ of Online Makespan Scheduling under Scenarios$(m,K)$ (OMSS$(m,K)$) is parametrized by a number $m\in\mathbb{N}$ of machines and a number $K\in\mathbb{N}$ of scenarios, and defined as a tuple consisting of a number $n\in\mathbb{N}$ of jobs, a processing time $p_{j}\geq 0$ for each job $j$ and job subsets $S_{1},\ldots,S_{K}\subseteq[n]$, called scenarios.
+Throughout this paper, for $n\in\mathbb{N}$, we denote by $n]\coloneqq\{1,\ldots,n\}$ the set of numbers $1$ through $n$.
+An instance $I=(n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ of Online Makespan Scheduling under Scenarios(m,K)(m,K) (OMSS(m,K)(m,K)) is parametrized by a number $m\in\mathbb{N}$ of machines and a number $K\in\mathbb{N}$ of scenarios, and defined as a tuple consisting of a number $n\in\mathbb{N}$ of jobs, a processing time $p_{j}\geq 0$ for each job $j$ and job subsets $S_{1},\ldots,S_{K}\subseteq[n]$, called scenarios.
 As we usually work with a single instance at a time, we often suppress this notation in the sequel as well as the parameters $m$ and $K$, which are often specified per section.
@@
 Our objective is to find a partitioning $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ of the set of jobs that minimizes $\max_{k\in[K]}\max_{i\in[m]}p(J_{i}\cap S_{k}).$ In other words, we consider schedules that arise from restricting a partition to each of the scenarios, and our goal is then to minimize the makespan of the worst-case scenario, which we also call the makespan of the schedule.
-Furthermore, our instance is online in the sense that parameters $m$ (number of machines) and $K$ (the number of scenarios) are known in advance, but jobs $j\in[n]$ as well as their processing times $p_{j}$ and the indices $k\in[K]$ such that $j\in S_{k}$ are revealed one job $j$ at a time.
-Once the information of a job $j$ is revealed in addition to already known information about the jobs in $[j-1]$, the job $j$ must be assigned to a machine $i\in[m]$, which is an irrevocable decision.
+Furthermore, our instance is online in the sense that the number $m$ of machines is known in advance, but jobs $j\in[n]$ as well as their processing times $p_{j}$ and the indices $k\in[K]$ such that $j\in S_{k}$ are revealed one job $j$ at a time.
+Once the information of a job $j$ is revealed in addition to already known information about the jobs in [j1][j-1], the job $j$ must be assigned to a machine $i\in[m]$, which is an irrevocable decision.
 In our analyses and arguments, we use the notion of a completion time of a job $j$.
@@
 Accordingly, we define the completion time of job $j\in J_{i}\cap S_{k}$ in scenario $S_{k}$ as $C^{k}_{j}\coloneqq\sum_{j^{\prime}\in J_{i}\cap S_{k}\cap[j]}p_{j^{\prime}}$ and the completion time as $C_{j}\coloneqq\max_{k\in[K],j\in S_{k}}C_{j}^{k}$.
+We can further define
+$\mathrm{makespan}(j)\coloneqq\max_{j^{\prime}\in[j]}C_{j^{\prime}}.$
 Online algorithms are often analyzed by their performance compared to the best possible outcome under full information.
@@
 In this special case, we may view our jobs as nodes, our scenarios as hyperedges containing a subset of the nodes and the partition $J_{1}\dot{\cup}\ldots\dot{\cup}J_{m}$ as a coloring $c\colon[n]\to[m]$ of the nodes.
-Then, our problem can be formulated as an online hypergraph coloring problem, in which we color a hypergraph $H=(V,E)$ with a fixed number of colors $1,\ldots,m$ and seek to minimize $\max_{i\in[m],e\in E}|\{c^{-1}(i)\}\cap e|$.
+Then, our problem can be formulated as an online hypergraph coloring problem, in which we color a hypergraph $\mathcal{H}=(V,E)$ with a fixed number of colors $1,\ldots,m$ and seek to minimize $\max_{i\in[m],e\in E}|\{c^{-1}(i)\}\cap e|$.
 The nodes $j\in[n]$ are revealed one at a time in increasing order along with partial hyperedges $e\cap[j]$ $(e\in E,j\in[n])$.
 [^1] According to this definition, hyperedges can have size $1$, although such hyperedges can and will be omitted as they pose no difference with respect to the objective function.
-We shall name this problem Online Makespan Hypergraph Coloring$(m)$ (OMHC$(m)$) and address it in Section 4.
+We shall name this problem Online Makespan Hypergraph Coloring(m)(m) and address it in Section 4.
+Our construction in Section 5 is even restricted to hyperforests.
+A hyperforest is a hypergraph without a hypercycle, where a hypercycle is defined as a sequence $v_{1},e_{1},v_{2}\ldots,e_{\ell},v_{\ell+1}=v_{1}$ with pairwise distinct hyperedges $e_{t}$ such that $v_{t},v_{t+1}\in e_{t}$ for all $t\in[\ell]$.
+Finally, it is worth mentioning some big numbers that we encounter in Section 5.
+In particular, we use Knuth’s arrow notation to denote $F(x)\coloneqq x\uparrow\uparrow 5\coloneqq x^{x^{x^{x^{x}}}}$ and the superlogarithm $\mathrm{slog}_{5}(x)\coloneqq F^{-1}(x)$.
 1.3 Related Work
@@
 Moreover, due to differences in their respective objective functions, this problem does not seem to have a direct relation to Online Makespan Hypergraph Coloring.
-However, our approach has implications for a variant of Online Hypergraph Coloring in which partial hyperedges are known (see Section 1.1).
+However, our approach conveniently enables implications for a variant of Online Hypergraph Coloring in which partial hyperedges are known (see Section 1.1).
 The reader may also notice similarities to the well-known discrepancy minimization problem (see e.g.
@@
 2 Extending Graham’s List Scheduling
-Our first observation is a simple algorithm that generalizes Graham’s List Scheduling Algorithm for OMSS$(m,K)$ and performs well when there are too many machines to force the schedule to create significant discrepancy.
+Our first observation is a simple algorithm that generalizes Graham’s List Scheduling Algorithm for OMSS(m,K)(m,K) and performs well when there are too many machines to force the schedule to create significant disbalance.
 Our algorithm beats the trivial bound of $m$ if $m>K+1$.
 We first observe that if the number $m$ of machines is large enough, we can apply the pigeonhole principle to find a machine that is not among the most loaded machines of any scenario.
-For $s\in[m-1]$ and $k\in[K]$, we call a machine $i$ $s$-favorable with respect to the $k$-th scenario if there are at least $s$ distinct machines $i^{\prime}\neq i$ that each have at least the load of machine $i$ in the $k$-th scenario, i.e., $p(J_{i^{\prime}}\cap S_{k})\geq p(J_{i}\cap S_{k})$.
+For $s\in[m-1]$ and $k\in[K]$, we call a machine $i$ $s$-favorable with respect to the $k$-th scenario if there are at least $s$ machines $i^{\prime}\neq i$ that have at least the load of machine $i$ in the $k$-th scenario, i.e., $p(J_{i^{\prime}}\cap S_{k})\geq p(J_{i}\cap S_{k})$.
 Observation 2.1.
@@
 The above algorithm is $\left(\frac{m-1}{\left\lceil\frac{m}{K}\right\rceil}+1\right)$-competitive for $m>K$.
-Notice that the aforementioned competitive ratio is strictly smaller than $K+1$.
-This algorithm is also applicable in the special case where every job appears in at most $k$ scenarios for a fixed $k\in\mathbb{N}$, yielding a ratio of $\nicefrac{{(m-1)}}{{\left\lceil\frac{m}{k}\right\rceil}}+1$.
-Furthermore, plugging in $K=1$ yields the well-known competitive ratio of $2-\nicefrac{{1}}{{m}}$ of Graham’s List Scheduling Algorithm.
+Proof 2.4.
+Let $i\in[m]$ be a machine and $k\in[K]$ a scenario that attains the makespan of the schedule, and let $j\in[n]$ be the smallest index such that $p(J_{i}\cap S_{k}\cap[j])=\mathrm{makespan}(n)$.
+We define $L\coloneqq p(J_{i}\cap S_{k}\cap[j-1])=\mathrm{makespan}(n)-p_{j}$.
+By the choice of $i$ upon the insertion of $j$, there are at least $\left\lceil\frac{m}{K}\right\rceil$ machines $i^{\prime}$ with $p(J_{i^{\prime}}\cap S_{k}\cap[j-1])\geq L$, including $i$, so that the offline optimum is bounded from below by
+$\operatorname{\mathrm{offl}}$ $\geq\max\left\{p_{j},\frac{p(S_{k})}{m}\right\}\geq\max\left\{p_{j},\frac{p(S_{k}\cap[j-1])+p_{j}}{m}\right\}$ $\geq\max\left\{p_{j},\frac{\left\lceil\frac{m}{K}\right\rceil\cdot L+p_{j}}{m}\right\}.$
+Hence, for $p_{j}\geq\frac{\left\lceil\frac{m}{K}\right\rceil}{m-1}L$, the competitive ratio $\rho$ is bounded by
+$\rho\leq\frac{L+p_{j}}{p_{j}}\leq 1+\frac{m-1}{\left\lceil\frac{m}{K}\right\rceil},$
+and otherwise it is also bounded by
+$\rho\leq\frac{L+p_{j}}{\frac{\left\lceil\frac{m}{K}\right\rceil\cdot L+p_{j}}{m}}=m\cdot\frac{L+p_{j}}{\left\lceil\frac{m}{K}\right\rceil\cdot L+p_{j}}\leq m\cdot\frac{1+\frac{\left\lceil\frac{m}{K}\right\rceil}{m-1}}{\left\lceil\frac{m}{K}\right\rceil+\frac{\left\lceil\frac{m}{K}\right\rceil}{m-1}}=1+\frac{m-1}{\left\lceil\frac{m}{K}\right\rceil}.$
+Here, the second inequality holds since $\left\lceil\frac{m}{K}\right\rceil>1$.
+Corollary 2.5.
+The above algorithm is $(K+1)$-competitive.
+This algorithm is also applicable in the special case where every job appears in at most $k$ scenarios for a fixed $k\in\mathbb{N}$, yielding a ratio of $\frac{m-1}{\left\lceil\frac{m}{k}\right\rceil}+1$.
+Furthermore, plugging in $K=1$ yields the well-known competitive ratio of $2-\frac{1}{m}$ of Graham’s List Scheduling Algorithm.
 3 Near-Tight Bounds for $m=2$ Machines
@@
 To show this, we present a minimalistic yet insightful algorithm (Algorithm 1) for makespan scheduling under $K=2$ scenarios and $m=2$ machines.
+To this end, we combine ideas about what the algorithm can feature with observations about how a worst-case instance is bound to look under the ideas we have fixed.
+To this end, we first need to introduce a few notions and machinery.
+3.1 Preparation for the Algorithm
+Fixing a rule.
 Our general idea is to fix a simple rule for assigning double-scenario jobs, i.e., jobs in $S_{1}\cap S_{2}$, and partition the single-scenario jobs $j\in S_{1}\triangle S_{2}$ so that the machines are reasonably balanced even after upcoming hypothetical double-scenario jobs.
 Rule 3.3.
-If $j\in S_{1}\cap S_{2}$, then assign the job $j$ to a machine $i$ such that the makespan of the schedule was previously attained by the other machine $3-i$.
-In case of a tie, select the machine that received the job $j-1$ ($j=1$ is assigned to the first machine if applicable).
+If $j\in S_{1}\cap S_{2}$, then assign the job $j$ to a machine $i$ such that the makespan of the schedule was previously attained by the other machine 3i3-i.
+In case of a tie, select the machine that received the job j1j-1 ($j=1$ is assigned to the first machine if applicable).
+A more approachable notion of competitiveness.
 Throughout our analysis, we only use two types of lower bounds for the offline optimum: Processing times $p_{j}$ of selected jobs $j$ whose completion times equal the makespan, and the average load of a selected scenario $S_{k}$.
@@
 We define the proxy competitive ratio of a schedule as
-$\rho=\frac{\max_{i,k\in\{1,2\}}p(S_{k}\cap J_{i})}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k})}{2}\right\},\max_{j\in[n]\cap\mathrm{argmax_{j\in[n]}\{C_{j}\}}}\{p_{j}\}\right\}},$
+$\rho=\frac{\max_{i,k\in\{1,2\}}p(S_{k}\cap J_{i})}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k})}{2}\right\},\max_{j\in[n]\cap\mathrm{argmax\{C_{j}\}}}\{p_{j}\}\right\}},$
 where $C_{j}$ denotes the completion time of the job $j$.
 The proxy competitive ratio is an upper bound on the competitive ratio.
-The lower bounds that we use are already well-known; e.g., Graham’s List Scheduling Algorithm is analyzed by bounding the proxy competitive ratio from above.
-We take this approach, however, one step further and restrict ourselves to analyzing only some instances that attain the largest possible proxy competitive ratio (see next Observation).
-This is significantly more difficult for the competitive ratio in the usual sense, as a witness to having good or bad offline optima highly depends on how the processing times of jobs fit together numerically.
-Observation 3.4 (cf. Figure 1).
-Let $\mathcal{A}$ be an online algorithm for OMSS(2,2)22(2,2)( 2 , 2 ) obeying Rule 3.3. Assume that $\mathcal{A}$ has a proxy competitive ratio of at most $\frac{5}{3}$ on instances where
+The lower bounds that we use are already well-known; for instance, Graham’s List Scheduling Algorithm is in fact analyzed by bounding the proxy competitive ratio from above.
+We take this approach, however, one step further and restrict ourselves to analyzing only the instances that attain the largest possible proxy competitive ratio.
+This would have been significantly more difficult for the competitive ratio in the usual sense, as a witness to having good or bad offline optima highly depends on how the processing times of jobs fit together numerically.
+Transforming to a worst-case instance.
+In search of instances with the worst possible proxy competitive ratio, we transform arbitrary instances to those that admit a certain structure, and are no better in terms of competitiveness, through a few simple operations.
+Definition 3.
+Let $\mathcal{I}=((n,(p_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]}))$ be an instance of OMSS(2,2)(2,2).
+The deletion of a job $j$ is the sole modification that $p_{j}$ is set to 0, i.e., it gives rise to an instance $\mathcal{I}^{\prime}=(n,(p^{\prime}_{j})_{j\in[n]},\{S_{k}\}_{k\in[K]})$ with
+$p^{\prime}_{j^{\prime}}=\begin{cases}0,&j=j^{\prime}\\ p_{j},&j\neq j^{\prime}.\end{cases}$
+Lemma 3.4 (Deletion lemma, cf. Figure 1).
+Let $\mathcal{I}$ be an instance on which an algorithm $\mathcal{A}$ admits a competitive ratio $\rho$.
+Let $j\in[n]$ be the largest index such that $j\in S_{1}\triangle S_{2}$ holds.
+Assume furthermore that the assignment of the $j$-th job to the $i=\tau(j)$-th machine was executed so that the scenario more represented in the machine did not change after the $j$-th job, i.e.,
+$\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j-1])=\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j]).$
+Then the instance $\mathcal{I}^{\prime}$ obtained by the deletion of $j$ admits a (proxy) competitive ratio $\rho^{\prime}\geq\rho.$
+Proof 3.5.
+Let $\tau,\tau^{\prime}\colon[n]\to\{1,2\}$ denote the assignment of the original instance and and the modified instance, respectively.
+Both instances admit the identical assignment up to (and including) job j1j-1.
+We observe that
+$\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j-1])=\mathrm{argmax}_{k\in\{1,2\}}p(J_{i}\cap S_{k})$
+for $i\in\{1,2\}$ for $\mathcal{I}$ and analogously for $\mathcal{I}^{\prime}$, since $\{j+1,\ldots,n\}\subseteq S_{1}\cap S_{2}$ by the choice of $j$ and therefore the difference $p(J_{i}\cap S_{1})-p(J_{1}\cap S_{2})$ stays constant after the assignment of the $j$-th job.
+For both instances, the jobs $j+1,\ldots,n$ are distributed to a machine $i$ such that the other machine 3i3-i admits makespan by the same reasoning.
+Since the assignment of the job $j$ does not affect $\max_{k\in\{1,2\}}p(J_{i}\cap S_{k}\cap[j^{\prime}])$ for any $j^{\prime}\in[n]$, we have $\tau(j^{\prime})=\tau^{\prime}(j^{\prime})$ for $j^{\prime}>j$.
+The $j$-th job in instance $\mathcal{I}$ cannot be in a combination of machine and scenario that admits the makespan of the schedule.
+Thus the makespan $\operatorname{\mathrm{onl}}_{\mathcal{I}}$ and $\operatorname{\mathrm{onl}}_{\mathcal{I}^{\prime}}$ of both instances coincide.
+On the other hand, the offline optimal value or the denominator of the proxy competitive ratio do not increase by the modification, so that $\rho^{\prime}\geq\rho$ holds.
+[…]
+Figure 1: Deletion of the largest $j\in S_{1}\triangle S_{2}$ (checkered) does not increase the approximation ratio, as it does not affect the execution of Rule 3.3.
+We may assume for the completion time of the largest job index $j\in S_{1}\triangle S_{2}$ that $C_{j}>C_{j-1}$, for otherwise, we may apply Lemma 3.4 successively to eliminate jobs $j\in S_{1}\triangle S_{2}$ that cause a violation of this inequality and analyze the resulting instances instead.
+Next, we observe that under certain circumstances, we may “cut” the double-jobs $j$ in that we replace them by other jobs $j^{\prime},j^{\prime\prime}\in S_{1}\cap S_{2}$ such that $p_{j^{\prime}}+p_{j^{\prime\prime}}=p_{j}$, while ensuring that the behavior of our algorithm remains the same way:
+Lemma 3.6 (Cutting lemma, cf. Figure 2).
+Let $t\in S_{1}\cap S_{2}$ be a job index for an instance $\mathcal{I}=(n,(p_{j})_{j\in[n]},\{S^{\prime}_{k}\}_{k\in\{1,2\}})$ of OMSS(2,2) such that $\mathrm{makespan}(t)>\mathrm{makespan}(t-1)$.
+Consider a schedule $\tau\colon[n]\to\{1,2\}$ of $n$ jobs, on which the algorithm admits a (proxy) competitive ratio of $\rho$.
+Then there is another instance $\mathcal{I}^{\prime}=(n+1,(p^{\prime}_{j})_{j\in[n]},\{S^{\prime}_{k}\}_{k\in\{1,2\}})$ with $n+1$ jobs with processing times $p^{\prime}_{j},j\in[n+1]$ under scenarios $S_{1}^{\prime},S_{2}^{\prime}\subseteq[n+1]$ with the following properties:
+  1. (i)
+    $p^{\prime}(J^{\prime}_{i}\cap S^{\prime}_{k}\cap[j])=p(J_{i}\cap S_{k}\cap[j])$ for $j\in[t-1]$, $i\in\{1,2\}$ and $k\in\{1,2\}$.
+  2. (ii)
+    The algorithm assigns the first $t$ jobs of $\mathcal{I}^{\prime}$ so that
+    $\max_{k\in\{1,2\}}p(J^{\prime}_{1}\cap S^{\prime}_{k}\cap[t])=\max_{k\in\{1,2\}}p(J^{\prime}_{2}\cap S^{\prime}_{k}\cap[t]).$
+  3. (iii)
+    The algorithm admits a (proxy) competitive ratio $\rho^{\prime}\geq\rho$ on the instance $\mathcal{I}^{\prime}$.
+Proof 3.7.
+We define the new instance $\mathcal{I}^{\prime}$ as follows: we set
+$S^{\prime}_{k}=(S_{k}\cap[t])\;\dot{\cup}\;[t+1]\;\dot{\cup}\;\{j\in\{t+2,\ldots,n+1\}\colon j-1\in S_{k}\}$
+for $t\in[n]$ and $k\in\{1,2\}$.
+Further, we define $p^{\prime}_{j}=p_{j}$ for $1\leq j<t$ and $p^{\prime}_{j}=p_{j-1}$ for $t+2\leq j\leq n+1$.
+We determine $p^{\prime}_{t}$ and $p^{\prime}_{t+1}$ later.
+We will aim to satisfy $p^{\prime}_{t}+p^{\prime}_{t+1}=p_{t}$.
+We start by noticing that the instances $\mathcal{I}$ and $\mathcal{I}^{\prime}$ coincide when restricted to the first t1t-1 jobs.
+Therefore, the algorithm will yield identical assignments until the assignment of the $t$-th job and the first property is satisfied.
+Now we consider the job indexed $t$.
+In both instances, we have $t\in S_{1}\cap S_{2}$ resp.
+$t\in S_{1}^{\prime}\cap S_{2}^{\prime}$.
+For such a double-scenario job, the assignment only depends on the prior subschedule $\tau|_{[t-1]}=\tau^{\prime}|_{[t-1]}$ and not on the processing time $p_{t}$ resp.
+$p_{t}^{\prime}$.
+Therefore, we also have $\tau(t)=\tau(t^{\prime})$ for any $p^{\prime}_{t}\in(0,p_{t})$.
+To obtain the second property, we pick
+$p^{\prime}_{t}=\mathrm{makespan}(t-1)-\min_{i^{\prime}\in\{1,2\}}\max_{k^{\prime}\in\{1,2\}}p(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t-1]),$ (1)
+and $p^{\prime}_{t+1}=p_{t}-p^{\prime}_{t}$.
+We guarantee that $p^{\prime}_{t+1}>0$ because of $\mathrm{makespan}(t)>\mathrm{makespan}(t-1)$.
+Since the $t$-th job is assigned to the $i^{\prime}$-th machine attaining the minimum in (1), we then have
+$\max_{k\in\{1,2\}}p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t])=\max_{k\in\{1,2\}}p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t-1])+p^{\prime}_{t}=\mathrm{makespan}(t-1).$ (2)
+Since the $t$-th job was not assigned to the $(3-i^{\prime})$-th machine, we also have
+$\max_{k\in\{1,2\}}\{p^{\prime}(J^{\prime}_{3-i^{\prime}}\cap S^{\prime}_{k}\cap[t])\}=\max_{k\in\{1,2\}}\{p^{\prime}(J^{\prime}_{3-i^{\prime}}\cap S^{\prime}_{k}\cap[t-1])\}=\mathrm{makespan}(t-1).$ (3)
+Comparing lines (2) and (3) reveals that the second property is satisfied.
+For the third property, we observe that both machines admit makespan prior to the insertion of the $(t+1)$-st job, so the algorithm will assign this job to the machine where it has assigned the previous job $t$ to.
+This yields
+$p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t+1])$ $=p^{\prime}(J^{\prime}_{i^{\prime}}\cap S^{\prime}_{k}\cap[t-1])+p^{\prime}_{t}+p^{\prime}_{t+1}$
+$=p(J_{i^{\prime}}\cap S_{k}\cap[t-1])+p_{t}=p(J_{i^{\prime}}\cap S_{k}\cap[t])$
+for $k\in\{1,2\}$.
+Since the $j$-th job in $\mathcal{I}$ is identical to the $(j+1)$-st job in $\mathcal{I}^{\prime}$, the algorithm will output the same value in both cases.
+It is also straightforward to observe that the denominator of the proxy competitive ratio cannot have increased after the modification.
+In total, this implies $\rho^{\prime}\geq\rho,$ proving that the third property is satisfied.
+[…]
+Figure 2: Execution of the cutting lemma on job $n$, before (left) and after (right).
+We observe that in order to analyze the proxy competitive ratio of instances whose last job $n$ satisfies $n\in S_{1}\cap S_{2}$, it suffices to consider those instances, in which additionally the first n1n-1 jobs satisfy
+$p(S_{k}\cap J_{k}\cap[n-1])=p(S_{3-k}\cap J_{3-k}\cap[n-1]).$ (4)
+Indeed, if we have an arbitrary instance with last job $n\in S_{1}\cap S_{2}$, we can apply Lemma 3.6 with $t=n$.
+On the other hand, if (4) is satisfied, the adversarial last job $n$ is double-scenario and has processing time
+$p_{n}=p(S_{k}\cap J_{k}\cap[n-1])+\max_{k\in\{1,2\}}p(S_{k}\cap J_{3-k}\cap[n-1])=\max_{k\in\{1,2\}}p(S_{k}\cap[n-1])$
+, so as to maximize the proxy competitive ratio
+$\rho$ $=\frac{p(S_{k}\cap J_{k}\cap[n-1])+p_{n}}{\max\left\{\max_{k\in\{1,2\}}\left\{\frac{p(S_{k}\cap[n-1])+p_{n}}{2}\right\},p_{n}\right\}}$ (5)
+$=\frac{2p(S_{k}\cap J_{k}\cap[n-1])+\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}{p(S_{k}\cap J_{k}\cap[n-1])+\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}$ (6)
+$=1+\frac{p(S_{k}\cap J_{k}\cap[n-1])}{\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}.$ (7)
+Now let $j\in[n]$ be the highest-indexed job that is in $S_{1}\triangle S_{2}$.
+If $j=n$, we assume to be done, so let $j<n$.
+By Lemma 3.4, there is a job $j^{\prime}>j$ whose completion time is larger than that of $j$ in some scenario.
+We can pick the smallest-indexed $j^{\prime}$ whose completion time in some scenario $S_{k}$ surpasses the completion time of $j$.
+For this index $j^{\prime}$, we apply Lemma 3.6 so that after the assignment of the earlier of the two jobs that replace $j^{\prime}$, the completion time of the job $j$ still equals the overall makespan of the schedule.
+This implies:
+Observation 3.8.
+Among the instances $\mathcal{I}=(n,(p_{j})_{j\in[n]},\{S_{1},S_{2}\})$ with $n\in S_{1}\cap S_{2}$, there is a worst-case instance with the following properties:
+  1. (i)
+    If $j\in[n]$ is the largest index with $j\in S_{1}\triangle S_{2}$, there is another job indexed $j^{\prime}>j$ such that $C_{j}=C_{j^{\prime}}$.
+  2. (ii)
+    Equation (4) holds, i.e., $p(S_{k}\cap J_{k}\cap[n-1])=p(S_{3-k}\cap J_{3-k}\cap[n-1])$.
+We transform this instance further by deleting the jobs $j^{\prime}+1,\ldots,n-1$ and changing $p_{n}$ to $p(S_{k}\cap J_{k}\cap[j^{\prime}])+\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])$.
+Since $j^{\prime}$ is the penultimate job after the modification, the equation (4) holds for the modified instance as well.
+If $\rho^{\prime}$ denotes the new proxy competitive ratio after the modifications, then we have
+$\rho$ $=1+\frac{p(S_{k}\cap J_{k}\cap[n-1])}{\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[n-1])}$ (8)
+$=1+\frac{p(S_{k}\cap J_{k}\cap[j^{\prime}])+p(J_{k}\cap S_{k}\cap\{j^{\prime}+1,\ldots,n-1\})}{\max_{k^{\prime}\in\{1,2\}}(p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])+p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap\{j^{\prime}+1\ldots,n-1\}))}$ (9)
+$=1+\frac{p(S_{k}\cap J_{k}\cap[j^{\prime}])+p(J_{k}\cap\{j^{\prime}+1,\ldots,n-1\})}{\max_{k^{\prime}\in\{1,2\}}(p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])+p(J_{3-k^{\prime}}\cap\{j^{\prime}+1\ldots,n-1\}))}$ (10)
+$\stackrel{{\scriptstyle(*)}}{{\leq}}1+\frac{p(S_{k}\cap J_{k}\cap[j^{\prime}]))}{\max_{k^{\prime}\in\{1,2\}}p(S_{k^{\prime}}\cap J_{3-k^{\prime}}\cap[j^{\prime}])}=\rho^{\prime}.$ (11)
+Here, the inequality ()(*) follows by Observation 3.8.
+Since $\rho$ is the proxy competitive ratio of a worst-case instance among those with $n\in S_{1}\cap S_{2}$, we conclude that the modified instance is a worst-case instance as well.
+We can therefore further strengthen the properties sketched in Observation 3.8 by assuming that the jobs before the penultimate job $j^{\prime}=n-1$ and after the last job $j\in S_{1}\triangle S_{2}$ had smaller completion time than the job $j$.
+In summary, we can observe the following statement.
+Observation 3.9 (cf. Figure 3).
+Let $\mathcal{A}$ be an online algorithm for OMSS(2,2)(2,2) obeying Rule 3.3. Assume that $\mathcal{A}$ has a proxy competitive ratio of at most $\frac{5}{3}$ on instances where
   1. (i)
@@
     $\mathcal{A}$ outputs a schedule with the following property: If $j\in[n]$ is the largest index with $j\in S_{1}\triangle S_{2}$, then $C_{j}=C_{n-1}$.
-Then, $\mathcal{A}$ is $\frac{5}{3}$-competitive.
-The above observation motivates the following parameter of significant value.
+Then, $\mathcal{A}$ is $\nicefrac{{5}}{{3}}$-competitive.
 […]
-Figure 1: Illustrative description of Observation 3.4.
-Processing times $p_{j}$ are given by the widths.
-Given the dark blue schedule, a worst outcome turns out to be the light blue job $n-1$ with a processing time $p_{n-1}$ such that $C_{n-1}=C_{j}$, followed by the light red job $n$ with $p_{n}=p(S_{1}\cap[n-1])$.
+Figure 3: Illustrative description of Observation 3.9.
+Given the dark blue schedule, a worst outcome turns out to be the light blue job n1n-1 with a processing time $p_{n-1}$ such that $C_{n-1}=C_{j}$, followed by the light red job $n$ with $p_{n}=p(S_{1}\cap[n-1])$.
+The anticipation measures the minimum discrepancy in either machine prior to the revelation of the final job $n$.
+With Lemma 3.4 in mind, we can conclude: among instances with $n\in S_{1}\cap S_{2}$, at least one worst-case instance behaves as follows: The last single-scenario job $j\in S_{1}\triangle S_{2}$ admits the makespan at the time of its insertion.
+Then, the double-scenario jobs $j+1,\ldots,n-1$ are all assigned to the opposite machine.
+After the insertion of the (n1)(n-1)-st job, both machines have the same makespan; i.e., the schedule satisfies (4).
+The last job $n$ has the adversarial processing time of
+$p_{n}=\max_{k\in\{1,2\}}p(S_{k}\cap[n-1])=p(S_{k}\cap J_{k}\cap[n-1])+\max_{k\in\{1,2\}}p(S_{k}\cap J_{3-k}\cap[n-1]).$
+Figure 3 illustrates a “worst” possible outcome, given a partial schedule.
+Anticipation.
+In the sequel, we assume that the instance which we analyze satisfies the properties in Observation 3.9.
+For such a special instance, a key parameter turns out to capture discrepancies within each machine.
 Definition 4.
 Consider a schedule where, without loss of generality, the makespan is attained at $S_{1}\cap J_{1}$. The anticipation $\alpha$ of this schedule is given by the ratio
-$\alpha\coloneqq\begin{cases}\frac{p(J_{1}\cap S_{1})}{\max\left\{p(J_{1}\cap S_{2}),p(J_{2}\cap S_{1})+p(J_{1}\cap S_{1})-p(J_{2}\cap S_{2})\right\}}&\text{if }p(J_{2}\cap S_{2})>p(J_{2}\cap S_{1}),\\ 0&\text{ else.}\end{cases}$
+$\alpha\coloneqq\begin{cases}\frac{p(J_{1}\cap S_{1})}{\max\left\{p(J_{1}\cap S_{2}),p(J_{2}\cap S_{1})+p(J_{1}\cap S_{1})-p(J_{2}\cap S_{2})\right\}}&\text{ if }p(J_{2}\cap S_{2})>p(J_{2}\cap S_{1})\\ 0&\text{ else.}\end{cases}$
 By convention, we assume $\frac{0}{0}=1$.
-In Figure 1, we have $p(J_{2}\cap S_{2}\cap[n-2])>p(J_{2}\cap S_{1}\cap[n-2])$ and the second term of the denominator corresponds to the load $p(J_{2}\cap S_{1}\cap[n-1])$, i.e., the load of $J_{2}\cap S_{1}$ after the $(n-1)$-st job is assigned.
-In general, anticipation measures the minimum discrepancy of a machine with respect to different scenarios after machine $2$ is loaded with double-scenario jobs just enough to attain the makespan of the schedule together with machine $1$.
-In our algorithm, one of the goals is to keep $\alpha\leq 2$ in order to brace ourselves for upcoming double-scenario jobs.
+In Figure 3, we have $p(J_{2}\cap S_{2}\cap[n-2])>p(J_{2}\cap S_{1}\cap[n-2])$ and the second term of the denominator corresponds to the load $p(J_{2}\cap S_{1}\cap[n-1])$, i.e., the load of $J_{2}\cap S_{1}$ after the (n1)(n-1)-st job is assigned.
+To further motivate the concept of anticipation, let us bound the proxy competitive ratio in terms of it.
+Lemma 3.10.
+Let a schedule of n1n-1 jobs be given that satisfies (4) and anticipation $\alpha$.
+After any assignment of the $n$-th job, the resulting schedule has proxy competitive ratio at most $\frac{2\alpha+1}{\alpha+1}$.
+Proof 3.11.
+Let $P\coloneqq p(S_{1}\cap J_{1}\cap[n-1])$. Then if $p_{n}\leq P+\frac{P}{\alpha}$, we have
+$\rho\leq\frac{2P+2p_{n}}{P+{\max\{p(S_{1}\cap J_{1}),p(S_{2}\cap J_{2})\}+p_{n}}}\leq\frac{2P+2p_{n}}{P+\frac{P}{\alpha}+P\left(1+\frac{1}{\alpha}\right)}\leq\frac{2\alpha+1}{\alpha+1},$
+and otherwise,
+$\rho\leq\frac{P+p_{n}}{p_{n}}\leq\frac{P+\left(1+\frac{1}{\alpha}\right)P}{\left(1+\frac{1}{\alpha}\right)P}=\frac{2\alpha+1}{\alpha+1},$
+showing the claim.
+3.2 The Algorithm
+In our algorithm, one of the goals is to keep $\alpha\leq 2$ in order to brace ourselves for upcoming double-scenario jobs in order to ensure a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$.
 The invariant which we maintain to this end is the following.
-Invariant 3.5.
+Invariant 3.12.
 The schedule satisfies the following assertions:
@@
 The last property might appear somewhat unnatural at the first glance, though it is crucial for bounding the anticipation $\alpha\leq 2$ in the subsequent iterations.
-Algorithm 1 shows how a single job $j$ is assigned given a schedule of the first $j-1$ jobs where Invariant 3.5 is maintained.
-To analyze the algorithm, we prove that instances satisfying the properties in Observation 3.4 are always assigned so that the first $n-1$ jobs maintain Invariant 3.5 and the last job results in a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$.
-Algorithm 1 The assignment of a single job $j$ given a $\frac{5}{3}$-competitive schedule that satisfies Invariant 3.5.
+Algorithm 1 shows how a single job $j$ is assigned given a schedule of the first j1j-1 jobs where Invariant 3.12 is maintained.
+In the sequel, we prove that instances satisfying the properties in Observation 3.9 are always assigned so that the first n1n-1 jobs maintain Invariant 3.12 and the last job results in a proxy competitive ratio of at most $\nicefrac{{5}}{{3}}$.
+Algorithm 1 The assignment of a single job $j$ given a $\frac{5}{3}$-competitive schedule that satisfies Invariant 3.12.
 1:Rename machines and scenarios so that $(J_{1}\cap S_{1})$ is maximal
@@
 5:if $j\in S_{2}\setminus S_{1}$ then
-6:     if assigning $j$ to the first machine satisfies Invariant 3.5 then
+6:  if assigning $j$ to the first machine satisfies Invariant 3.12 then
 7:   return 1
@@
 14:end if
-A significant portion of the analysis is focused on proving that if line 9 is executed, Invariant 3.5 is indeed maintained. We refer the reader to the full version of the paper for further details.
+3.3 Analysis of Algorithm 1
+It is evident that if Line 7 is executed, the assumption is maintained and we obtain a $\nicefrac{{5}}{{3}}$-competitive schedule.
+It is also easy to observe that the execution of Line 3 provides this: In the first scenario, the second machine is a least loaded machine, so that $\rho$ or $\alpha$ cannot increase.
+[^2] This is straightforward to observe and left to the reader as an exercise.
+Moreover, if the schedule was previously dominated by the first machine, the third property is maintained as we do not increase $p(J_{1}\cap S_{1})$.
+Finally, we deal with the execution of Line 13: the proxy competitive ratio does not exceed $\nicefrac{{5}}{{3}}$ after the last single-scenario job, which we prove below.
+Lemma 3.13.
+For an instance of OMSS(2,2) satisfying the properties in Observation 3.9, let $j\in[n]$ be the largest index with $j\in S_{1}\triangle S_{2}$.
+Then, if Algorithm 1 satisfies Invariant 3.12 on the first $j$ jobs, then it also satisfies Invariant 3.12 on the first $\ell$ jobs for $\ell\in\{j+1,\ldots,n-1\}$.
+Proof 3.14.
+  1. (i)
+    $\alpha\leq 2$: We claim that $\alpha$ stays constant after the insertion of the $j$-th job.
+Indeed, $p(J_{2}\cap S_{2}\cap[\ell])>p(J_{2}\cap S_{1}\cap[\ell])$ is maintained (as $p(J_{2}\cap S_{2}\cap[\ell])-p(J_{2}\cap S_{1}\cap[\ell])$ is constant).
+Similarly, both $p(J_{1}\cap S_{2}\cap[\ell])$ and $p(J_{2}\cap S_{1}\cap[\ell])+p(J_{1}\cap S_{1}\cap[\ell])-p(J_{2}\cap S_{2}\cap[\ell])$ are constant in $\ell\in\{j+1,\ldots,n\}$.
+  2. (ii)
+    $\rho\leq\nicefrac{{5}}{{3}}$: By Lemma 3.10, it suffices to observe that the anticipation prior to the assignment of the jobs $\ell\in\{j+1,\ldots,n\}$ is constant.
+  3. (iii)
+    Third property: If there was no machine dominance after the insertion of the $j$-th job, machine dominance does not emerge since the jobs $\ell\in\{j+1,\ldots,n-1\}$ are all assigned to the second machine.
+Otherwise, the condition of the third property is maintained for the same reason.
+This lemma does not cover the $n$-th job, however, the $n$-th job only needs to satisfy $\rho\leq\nicefrac{{5}}{{3}}$, which follows by the fact that $\alpha\leq 2$ prior to its assignment as we have seen, and Lemma 3.10.
+All that remains to analyze is that if Line 9 is executed, Invariant 3.12 is maintained.
+Notice that machine dominance cannot occur as the prior makespan is still among the top two largest loads and is in the opposite machine of the makespan.
+Hence the third property does not need to be checked.
+We apply a case distinction.
+In each case, we either bound the proxy competitive ratio $\rho\leq\nicefrac{{5}}{{3}}$ and the anticipation $\alpha\leq 2$ by means of inequalities, or find inequalities that are necessary for a condition to fail and show that conjunction of these conditions has no solutions.
+As all inequalities are linear in $p(J_{i}\cap S_{k}\cap[j-1])$ and $p_{j}$, the reader might interpret the remainder of the proof as a sequence of cheerful linear inequalities which they are encouraged to trace.
+Accordingly, we ease the notation, and define:
+$x_{1}$ $\coloneqq p(J_{1}\cap S_{1}\cap[j-1])$
+$x_{2}$ $\coloneqq p(J_{1}\cap S_{2}\cap[j-1])$
+$x_{3}$ $\coloneqq p(J_{2}\cap S_{1}\cap[j-1])$
+$x_{4}$ $\coloneqq p(J_{2}\cap S_{2}\cap[j-1])$
+$x_{5}$ $\coloneqq p_{j}$
+We start with the case that $j$ has a large processing time.
+Case 1: $x_{5}\geq 2x_{1}-x_{2}$.
+In this case, the attempt to assign the job $j$ to the first machine failed because
+$p(J_{1}\cap S_{2}\cap[j])=x_{2}+x_{5}\geq x_{2}+2x_{1}-x_{2}=2x_{1},$
+so that the assignment to the second machine yields a makespan of
+$p(J_{2}\cap S_{2}\cap[j])=x_{4}+x_{5}\geq x_{4}+2x_{1}-x_{2}\geq x_{1}.$
+The proxy competitive ratio is at most
+$\rho\leq\max\left\{\frac{x_{4}+x_{5}}{x_{5}},\frac{2x_{4}+2x_{5}}{x_{2}+x_{4}+x_{5}}\right\}\leq\frac{5}{3}.$
+Indeed, if $x_{1}\geq 2x_{2}$, then we have
+$\frac{x_{4}+x_{5}}{x_{5}}\leq\frac{x_{4}+2x_{1}-x_{2}}{2x_{1}-x_{2}}\leq\frac{3x_{1}-x_{2}}{2x_{1}-x_{2}}\leq\frac{2.5x_{1}}{1.5x_{1}}=\frac{5}{3},$
+which is equivalent to $2x_{5}\geq 3x_{4}$. Otherwise we have
+$\frac{2x_{4}+2x_{5}}{x_{2}+x_{4}+x_{5}}\leq\frac{5x_{4}}{x_{2}+2.5x_{4}}\leq\frac{5x_{1}}{0.5x_{1}+2.5x_{1}}\leq\frac{5}{3}.$
+Finally, the anticipation is bounded by
+$\alpha\leq\frac{x_{4}+x_{5}}{x_{2}+x_{4}+x_{5}-x_{1}}\leq\frac{x_{5}}{x_{2}+x_{5}-x_{1}}\leq\frac{2x_{1}-x_{2}}{x_{2}+2x_{1}-x_{2}-x_{1}}=\frac{2x_{1}-x_{2}}{x_{1}}\leq 2,$
+so that all properties in Invariant 3.12 are shown.
+In the sequel, we assume that $x_{5}\leq 2x_{1}-x_{2}$ and distinguish some cases based on how $x_{1},x_{2},x_{3},x_{4}$ relate to each other.
+Case 2: $x_{1}\geq x_{2}\geq x_{3}$ and $x_{2}\geq x_{4}$.
+The $j$-th job was assigned to the second machine.
+This cannot be because of properties (ii) or (iii), as these could only be violated if $x_{5}\geq 2x_{1}-x_{2}$, which we have already excluded.
+Hence, assigning the $j$-th job to the first machine must have created a proxy competitive ratio of larger than $\frac{5}{3}$, i.e.,
+$\frac{2x_{2}+2x_{5}}{x_{2}+x_{4}+x_{5}}>\frac{5}{3}\Leftrightarrow x_{2}+x_{5}>5x_{4}.$ (12)
+If the makespan is still attained by $J_{1}\cap S_{1}$, then the proxy competitive ratio $\rho$ has not increased and is bounded by $\frac{5}{3}$ by induction.
+Otherwise the new makespan is $x_{4}+x_{5}\geq x_{1}$ and the proxy ratio is bounded by $\rho\leq\frac{x_{4}+x_{5}}{\max\{x_{5},\frac{x_{2}+x_{4}+x_{5}}{2}\}}\leq\frac{3}{2}$, similar to the proof of Graham’s List Scheduling Algorithm since $j$ was assigned to a least loaded machine.
+We only need to check the property (ii), i.e., that $\alpha\leq 2$. However, the anticipation is bounded by $\frac{x_{1}}{x_{2}}\leq 2$ since the initial schedule must satisfy the property (iii).
+Case 3: $x_{1}\geq x_{3}$ and $x_{4}\geq x_{2}$.
+We claim that this case never occurs, that is, in this case the job $j$ is always assigned to the first machine.
+Since we assume $x_{5}\leq 2x_{1}-x_{2}$, the property (iii) cannot be the reason the assignment to the first machine has failed.
+Similarly, if $x_{5}+x_{2}\geq x_{1}$, the anticipation is bounded by $\frac{x_{5}+x_{2}}{x_{1}}\leq\frac{2x_{1}-x_{2}+x_{2}}{x_{1}}=2$, and otherwise the anticipation cannot increase, either; so the property (ii) is also maintained while assigning to the first machine.
+Finally, we note that by $x_{4}\geq x_{2}$, the first machine is least loaded in the second scenario, therefore by the proof of Graham’s List Scheduling, the proxy ratio either decreases or is bounded by $\frac{3}{2}$.
+Case 4: $x_{1}\geq x_{3}\geq x_{2}\geq x_{4}$, proof by LP solver.
+Exactly as above, the assignment cannot have failed because of the properties (ii) and (iii).
+Therefore, assigning $j$ to the first machine must yield a proxy competitive ratio larger than $\frac{5}{3}$, which implies both $2x_{5}\leq 3x_{2}$ and $x_{5}+x_{2}\leq 5x_{4}$.
+As $x_{4}\leq x_{2}$, the second machine is least loaded in the second scenario and therefore $\rho\leq\frac{5}{3}$ is secured if we place $j$ to the second machine.
+By the above observation, we can further bound the anticipation by
+$\alpha\leq\min\left\{\frac{x_{4}+x_{5}}{x_{3}+(x_{4}+x_{5}-x_{1})},\frac{x_{1}}{x_{2}+(x_{1}-x_{4}-x_{5})}\right\}.$
+If $x_{4}+x_{5}\geq x_{1}$, the first term is valid, which means
+$\frac{x_{4}+x_{5}}{x_{3}+(x_{4}+x_{5}-x_{1})}>2\Leftrightarrow 2(x_{1}-x_{3})>x_{4}+x_{5}.$ (13)
+Gathering all constraints, we obtain system of inequalities
+$x_{1}$ $\geq x_{3}$ (14)
+$x_{3}$ $\geq x_{2}$ (15)
+$x_{2}$ $\geq x_{4}$ (16)
+$2x_{5}$ $\leq 3x_{2}$ (17)
+$x_{4}+x_{5}$ $\geq x_{1}$ (18)
+$2x_{1}-2x_{2}$ $\geq x_{5}+x_{4}$ (19)
+$x_{4}+x_{5}$ $\geq 2x_{3}$ (20)
+$x_{2}+x_{5}$ $\geq 5x_{4}$ (21)
+The only feasible solutions to this system of inequalities are of the type $(4x_{4},2x_{4},2x_{4},x_{4},3x_{4})$.
+For these solutions, it is immediate that $\alpha\leq 2$, from which the claim follows.
+Similarly if $x_{4}+x_{5}\geq x_{1}$, the inequality (18) is reversed and the inequality (19) is replaced by $\frac{x_{1}}{x_{3}+(x_{1}-x_{4}-x_{5})}>2\Leftrightarrow 2x_{4}+2x_{5}\geq 2x_{3}+x_{1}$, so that all schedules to be inspected are the solutions of the following system:
+$x_{1}$ $\geq x_{3}$ (22)
+$x_{3}$ $\geq x_{2}$ (23)
+$x_{2}$ $\geq x_{4}$ (24)
+$2x_{5}$ $\leq 3x_{2}$ (25)
+$x_{4}+x_{5}$ $\leq x_{1}$ (26)
+$2x_{4}+2x_{5}$ $\geq 2x_{3}+x_{1}$ (27)
+$x_{4}+x_{5}$ $\geq 2x_{3}$ (28)
+$x_{2}+x_{5}$ $\geq 5x_{4}$ (29)
+This system of inequalities has the same set of solutions as well, so that the anticipation is bounded by $\alpha\leq 2$.
+As all cases are covered, we conclude that Algorithm 1 maintains the properties in Invariant 3.12 and is therefore $\nicefrac{{5}}{{3}}$-competitive.
+3.4 Lower Bounds
 In spite of the restrictions that arise from both the fixed double-scenario rule and how we evaluate lower bounds, we find out that our algorithm in the previous subsection is not far from being best possible.
@@
 For $m=2$ and $K=2$, there exists no deterministic algorithm with competitive ratio strictly smaller than $\frac{9+\sqrt{17}}{8}\approx 1.640$.
-Proof 3.7.
+Proof 3.16.
 To ease the notation once again, let $X\coloneqq 3+\sqrt{17}$.
+We demonstrate our family of instances by means of a case distinction, that is, our instances mostly agree with each other in earlier jobs and their later jobs are revealed depending on the schedule.
 The first four jobs are given by $1,2\in S_{1}\setminus S_{2}$, $3,4\in S_{2}\setminus S_{1}$, $p_{1}=p_{2}=p_{3}=p_{4}$.
@@
 In both cases, an eighth job $8\in S_{1}\cap S_{2}$ forces the schedules to the desired lower bound of competitiveness.
-Notice that the aforementioned family of instances fails to achieve small competitive ratios for reasons similar to those that require Property (iii) in Invariant 3.5.
-It can also be shown through a simple array of instances that, no algorithm that satisfies Rule 3.3 can beat the bound of $\nicefrac{{5}}{{3}}$, for which we refer the reader to the full version of the paper.
-4 Tight Bounds for Sufficiently Many Scenarios
-One of our main results (Theorem 1.2) states that for any number $m\in\mathbb{N}$ of machines, there exists a number $K$ of scenarios such that no deterministic algorithm for OMSS$(m,K)$ can have a competitive ratio less than the trivial upper bound $m$, even for unit processing times.
-Due to space constraints, we merely sketch a proof for the special case of $m=3$.
-While doing so, we showcase key ideas of the general argument, although we forfeit the restriction onto hypertrees for the sake of a more clearly structured description.
-In the context of scheduling, this is indeed not a significant change.
+One might wonder how the gap between the upper bound in Theorem 1.1 and the lower bound in Theorem 3.15 can be narrowed down.
+We can observe through a simple family of instances that, if that is possible at all by means of improving the upper bound, then dropping Rule 3.3 is necessary.
+Observation 3.17.
+Let $r<\nicefrac{{5}}{{3}}$ be arbitrary. There exists no deterministic $r$-competitive algorithm for OMSS(2,2) which satisfies Rule 3.3.
+Proof 3.18.
+Consider an instance of $n=5$ jobs with processing times $p_{1}=p_{2}=p_{4}=1$, $p_{3}=2$, $p_{5}=3$. The scenarios are given by $S_{1}=\{1,2,4,5\}$, $S_{2}=\{3,4,5\}$.
+The first two jobs must be assigned to opposite machines; for otherwise one could truncate the instance, i.e., consider the restricted instance of the first $n=2$ jobs with $p_{j}\equiv 1$, $S_{1}=\{1,2\}$, $S_{2}=\emptyset$ instead.
+On this restricted instance, the algorithm would be no better than $2$-competitive.
+Hence, we may assume without loss of generality that $1\in J_{1}$ and $2\in J_{2}$.
+The third job is assigned to the first machine without loss of generality.
+Then, as $4\in S_{1}\cap S_{2}$ and the first machine uniquely admits makespan, we must assign the fourth job to the second machine, and similarly, the fifth job is assigned to the second machine as well.
+In total, the makespan is given by
+$p(J_{1}\cap S_{2})=1+1+3=5,$
+although an optimal solution is given by $J_{1}=\{1,2,3,4\}$ and $J_{2}=\{5\}$, which has a makespan of $3$.
+[…]
+Figure 5: The online schedule in the proof of Observation 3.17 and the offline optimum.
+In the above setting, the right course of action would have indeed been to violate Rule 3.3 and assign $4\in J_{1}$.
+Hence, it is justified to suspect that the algorithm can be improved by altering this rule according to the sizes of jobs $j\in S_{1}\cap S_{2}$, although as seen in Theorem 3.15, it can only improve the ratio by less than $0.027$.
+4 Tight Bounds for $m=3$ and Sufficiently Many Scenarios
+One of our main results (Theorem 1.2) states that for any number $m\in\mathbb{N}$ of machines, there exists a number $K$ of scenarios such that no deterministic algorithm for OMSS(m,K)(m,K) can have a competitive ratio less than $m$, even for unit processing times.
 Theorem 4.1.
-For OMHC$(3)$, there exists no $r$-competitive online algorithm for any $r<3$, even when restricted to instances with at most $233$ hyperedges of size at most $3$ each.
-The problem OMHC$(m)$, which we formulated in Section 1.2, is equivalent to the unit processing time case of OMSS$(m,K)$ with $K$ part of the input, so that the lower bounds are immediately implied for the latter as well.
+Let $K\geq 233$.
+Then for Online Makespan Hypergraph Coloring(m)(m), there exists no $r$-competitive online algorithm for any $r<3$, even when the problem is restricted to $3$-uniform instances with at most $K$ hyperedges.
+The problem Online Makespan Hypergraph Coloring(m)(m), which we formulated in Section 1.2, is equivalent to the unit processing time case of OMSS(m,K)(m,K) with $K$ part of the input, so that the lower bounds are immediately implied for the latter as well for $K\geq 233$.
 To prove Theorem 4.1, we provide a family of instances on which any online algorithm is forced to output a monochromatic hyperedge of size $3$.
-For each of the instances, we also maintain (offline) colorings $c\colon[j]\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$, with the property that if two nodes $v,v^{\prime}$ are contained in a common hyperedge, we have $c(v)\neq c(v^{\prime})$.
-The latter coloring suggests an offline optimal value of $1$ for each of the instances, implying the competitive ratio of $3$ immediately.
-The main difference between the online and offline colorings is, as expected, that we may change the offline coloring completely at all times as long as its makespan equals $1$.
-Among other modifications, we often employ a permutation $\pi\colon\{\mathrm{red,blue,yellow}\}\to\{\mathrm{red,blue,yellow}\}$ on connected components to realize this possibility.
+For each of the instances, we also maintain (offline) colorings $c\colon[j]\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ to be feasible, i.e., with the property that if two nodes $v,v^{\prime}$ are included in a common hyperedge $e$, we have $c(v)\neq c(v^{\prime})$.
+The latter coloring suggests an offline optimum of value $1$ for each of the instances, implying the competitive ratio of $3$ immediately.
+The main difference between the online and offline colorings is, as expected, that we may change the offline coloring completely at all times as long as it remains feasible.
+Among other modifications, we often employ a permutation $\pi\colon\{\mathrm{red,blue,yellow}\}\to\{\mathrm{red,blue,yellow}\}$ to realize this possibility.
 We must trace the offline coloring simultaneously because the hyperedges are often defined adversarially according to both colorings of the current hypergraph.
-To differentiate between the online and offline colorings, we denote by $J_{i}\subseteq V$ the subset of nodes assigned to the online color $i$ for $i\in\{1,2,3\}$.
-Right before the revelation of a node $j$, we implicitly mean $J_{i}\cap[j-1]$ whenever we mention $J_{i}$ for some $i\in\{1,2,3\}$.
+In the following, let $J_{i}\subseteq V$ denote the subset of nodes assigned to the online color $i$ for $i\in\{1,2,3\}$.
+Right before the insertion of a node $j$, we implicitly mean $J_{i}\cap[j-1]$ whenever we mention $J_{i}$ for some $i\in\{1,2,3\}$.
 Inspired by our original scheduling problem, we call the objective value $\max_{e\in E}\max_{i\in\{1,2,3\}}|e\cap J_{i}|$ of an instance its makespan.
-Without loss of generality, we truncate an instance as soon as the online solution is forced to obtain a makespan of $3$.
-Our gadget: Seven-node subhypergraph $\mathcal{S}$.
-Initially, we study a subhypergraph $\mathcal{S}$ that we use as a building block throughout our construction.
+Without loss of generality, we can truncate an instance as soon as it obtains a makespan of $3$.
+4.1 Our gadget: Seven-node subgraph $\mathcal{S}$
+Initially, we study a subgraph $\mathcal{S}$ that we use as a building block throughout our construction.
 The hypergraph $\mathcal{S}$ is given by the nodes $V(\mathcal{S})=\{1,\ldots,7\}$ and hyperedges
+$E(\mathcal{S})=\big\{$ $\{1,2\},\{1,3\},\{2,3\},\{1,2,3\},\{2,4\},\{4,5\},\{2,5\},\{2,4,5\},$
+$\{4,6\},\{6,7\},\{4,7\},\{4,6,7\},\{3,6\},\{3,7\}\big\}.$
+One can verify that the hyperedges have size at most $3$ and the hyperedge set $E(\mathcal{S})$ is closed under taking subsets of size at least $2$, i.e., if $e\in E(\mathcal{S})$, $|e|=3$ and $e^{\prime}\subseteq e$ with $|e^{\prime}|=2$, then $e^{\prime}\in E(\mathcal{S})$.
+The hypergraph $\mathcal{S}$ together with the maximal hyperedges $e_{1},\ldots,e_{5}$ is given in Figure 6.
+$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$33112255446677
+Figure 6: The hypergraph $\mathcal{S}$ with the maximal hyperedges $e_{1},\ldots,e_{5}$.
+We first assume that the seven nodes are assigned to exactly two out of three online colors. Since the online colors are initially symmetric, let us assume that we use the online colors $1$ and $2$.
+Assumption 4.2.
+$e_{k}\cap J_{3}=\emptyset$ for all $k\in\{1,2,3\}$.
+Moreover, for brevity, we only mention the revelation of inclusion-wise maximal hyperedges whenever we reveal a copy of this gadget, although the subsets of these hyperedges of size $2$ are being introduced as well.
+First, we describe a case distinction about the introduction of the gadget, always satisfying Assumption 4.2.
+We reveal the first three nodes that appear in the common hyperedge $e_{1}=\{1,2,3\}$.
+If all three nodes were assigned to a common online color, we would obtain a makespan of $3$, which ultimately proves the theorem.
+Therefore, we assume that the nodes were partitioned in two online colors; as the nodes are symmetric so far, we may assume without loss of generality that the first two nodes were assigned to the first color and the third job to the second, see Figure 7.
 […]
+Figure 7: First three nodes.
+Now we introduce two more nodes $4,5$ in this order one by one and reveal that
+$e_{2}=\{2,4,5\}.$
+If both nodes $4$ and $5$ are assigned to the second online color, we may assign the remaining two nodes $6$ and $7$ arbtirarily and ignore them from now on, see Figure 8.
 […]
-The hyperedges have size at most $3$ and the set $E(\mathcal{S})$ is closed under taking subsets of size at least $2$, i.e., if $e\in E(\mathcal{S})$, $|e|=3$ and $e^{\prime}\subseteq e$ with $|e^{\prime}|=2$, then $e^{\prime}\in E(\mathcal{S})$.
-The hypergraph $\mathcal{S}$ together with the maximal hyperedges $e_{1},\ldots,e_{5}$ is given in Figure 3.
-$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$3333111122225555444466667777
-Figure 3: The hypergraph $\mathcal{S}$ with the maximal hyperedges $e_{1},\ldots,e_{5}$.
-We first assume that the seven nodes in $V(\mathcal{S})$ are assigned to exactly two out of the three online colors.
-Since the online colors are initially symmetric, let us assume that we use the online colors $1$ and $2$.
-By a careful case distinction, we may observe:
+Figure 8: The five relevant nodes if $\tau(4)=\tau(5)=2$.
+Otherwise, one of the nodes $4$ and $5$, without loss of generality, the node $4$, is assigned to the first color. We reveal the sixth node along with the hyperedges
+$e_{3}\supseteq\{4,6\},e_{4}\supseteq\{3,6\}.$
+If the sixth node is assigned to the second color, we may ignore the seventh node similar to above, see Figure 9.
+[…]
+Figure 9: The six relevant nodes if $\tau(6)=2$.
+If the sixth node is assigned to the first online color, we reveal the seventh node along with the hyperedges
+$e_{3}=\{4,6,7\},e_{5}\supseteq\{3,7\}.$
+Throughout the remainder of this section, we let $\mathcal{S}$ denote the resulting hypergraph with seven nodes and fourteen hyperedges (including subsets).
+Since both the sixth and the seventh nodes are assigned to the first online color, the seventh node is forced to be assigned to the second online color. Our final assignment looks as in Figure 10.
+[…]
+Figure 10: The hypergraph $\mathcal{S}$ together with the online and offline colorings, under the assumption that $\tau(6)=1$.
 Proposition 5.
-For any partitioning $V(\mathcal{S})=J_{1}\dot{\cup}J_{2}$ of the nodes that admits a makespan of at most $2$, there exists a node coloring $c\colon V(\mathcal{S})\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ such that there are two hyperedges $e\in E(\mathcal{S}[J_{1}]),e^{\prime}\in E(\mathcal{S}[J_{2}])$ with $c(v)\in\{\mathrm{red},\mathrm{blue}\}$ (or another subset of $\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ of size $2$) for all $v\in e\cup e^{\prime}$.
-In other words, we have induced monochromatic hyperedges of size $2$ with two different colors, both with respect to the online coloring.
-Applying this repeatedly, we can further force the algorithm to a monochromatic hyperedge of size $3$, which we show next.
-Pigeonholing the two-color case.
+For any partitioning $V(\mathcal{S})=J_{1}\dot{\cup}J_{2}$ of the nodes that admits a makespan of at most $2$, there exists a node coloring $c\colon V(\mathcal{S})\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ such that there are two hyperedges $e\in E(\mathcal{S}[J_{1}]),e^{\prime}\in E(\mathcal{S}[J_{2}])$ with $c(e)=c(e^{\prime})=\{\mathrm{red},\mathrm{blue}\}$.
+Proof 4.3.
+See Figures 8, 9 and 10 for the offline coloring in the respective cases.
+The hyperedges $(e,e^{\prime})$ in question are $(\{1,2\},\{4,5\})$; $(\{2,4\},\{3,6\})$ and $(\{2,4\},\{3,7\})$, respectively.
+In other words, we have induced monochromatic hyperedges of both colors with respect to the online coloring.
+Applying this repeatedly, we can force the algorithm to a monochromatic hyperedge of size $3$, which we show next.
+4.2 Pigeonholing the two-color case
 In our construction, we would like to create copies of the graph $\mathcal{S}$ successively, which we denote by $\mathcal{S}(t)$ for incrementing $t\in\mathbb{N}$.
-A straightforward application of the pigeonhole principle yields that, if we have seven subhypergraphs $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ that are each partitioned onto two online colors, at least three of them (without loss of generality $\mathcal{S}(1)$, $\mathcal{S}(2)$, $\mathcal{S}(3)$) must be partitioned onto the same two online colors.
+A straightforward application of the pigeonhole principle yields that, if we have seven subgraphs $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ that are each partitioned onto two online colors, at least three of them (without loss of generality $\mathcal{S}(1)$, $\mathcal{S}(2)$, $\mathcal{S}(3)$) must be partitioned onto the same two online colors.
 Again without loss of generality, these two online colors are $1$ and $2$.
-Considering Proposition 4, we can permute the colorings of the three copies $\mathcal{S}(1)$, $\mathcal{S}(2)$ and $\mathcal{S}(3)$ each offline so that for any offline color $u\in\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$, there exists a copy $\mathcal{S}(t)$ with two (online) monochromatic hyperedges $e_{i}^{t}\in E(\mathcal{S}(t))$ $(i\in\{1,2\})$ of size $2$ avoiding the color $u$.
-The hypergraph in Figure 4 restricted to nodes of the form $j_{s}^{i}$ illustrate these edges.
-Once the six edges in question are found, one can insert nodes $x_{1}$, $x_{2}$, $x_{3}$ in this order, revealing hyperedges as depicted in Figure 4.
-To avoid a monochromatic edge, both $x_{1}$ and $x_{2}$ must be assigned the third online color.
-Keeping this in mind, it follows that $x_{3}$ cannot be assigned an online color without creating a monochromatic edge with respect to the online coloring.
-Together with the fact that the offline coloring indeed has makespan $1$, we obtain:
+Considering Proposition 5, we can color the three copies $\mathcal{S}(1)$, $\mathcal{S}(2)$ and $\mathcal{S}(3)$ so as to satisfy the following properties:
+  • There exist two hyperedges $e_{i}^{1}=\{j_{i}^{1},j_{i}^{2}\}\in E(\mathcal{S}(1))$ for $i\in\{1,2\}$ with the property that $c(e_{i}^{1})=\{\mathrm{red,blue}\}$,
+  • there exist two hyperedges $e_{i}^{2}=\{j_{i}^{3},j_{i}^{4}\}\in E(\mathcal{S}(2))$ for $i\in\{1,2\}$ with the property that $c(e_{i}^{1})=\{\mathrm{blue,yellow}\}$,
+  • there exist two hyperedges $e_{i}^{3}=\{j_{i}^{5},j_{i}^{6}\}\in E(\mathcal{S}(3))$ for $i\in\{1,2\}$ with the property that $c(e_{i}^{3})=\{\mathrm{yellow,red}\}$.
+Moving forward, we restrict our subgraph to one with the nodes $j_{i}^{t}$ for $i\in\{1,2\},t\in\{1,\ldots,6\}$ and the hyperedges $k_{i}^{t}$ for $i\in\{1,2\},t\in\{1,2,3\}$.
+The remaining nodes in $V(\mathcal{S}(1))\dot{\cup}V(\mathcal{S}(2))\dot{\cup}V(\mathcal{S}(3))$ will not be connected to other nodes anymore, nor will their offline coloring be altered.
+Thus, they can be ignored from now on.
+$j_{2}^{1}$$j_{1}^{1}$$j_{2}^{2}$$j_{1}^{2}$$e_{1}^{1}$$e_{2}^{1}$$j_{2}^{3}$$j_{1}^{3}$$j_{2}^{4}$$j_{1}^{4}$$e_{1}^{2}$$e_{2}^{2}$$j_{2}^{5}$$j_{1}^{5}$$j_{2}^{6}$$j_{1}^{6}$$e_{1}^{3}$$e_{2}^{3}$$J_{1}$$J_{2}$
+Figure 11: The nodes $j_{i}^{t}$.
+Now we introduce the thirteenth node $x_{1}$, revealing the full extent of the hyperedges $e_{1}^{1}=\{j_{1}^{1},j_{1}^{2},x_{1}\}$ and $e_{2}^{1}=\{j_{2}^{1},j_{2}^{2},x_{1}\}$.
+Unless a competitive ratio of $3$ is attained, the node $x_{1}$ must be assigned to the third online color.
+Thereafter, the fourteenth node $x_{2}$ is revealed together with hyperedges $e_{1}^{2}=\{j_{1}^{3},j_{1}^{4},x_{2}\}$, $e_{2}^{2}=\{j_{2}^{3},j_{2}^{4},x_{2}\}$, $f_{1}\supseteq\{x_{1},x_{2}\}$.
+Analogously, the fourteenth node is assigned to the third online color as well.
+Notice that we can feasibly extend the offline coloring by $c(x_{1})=\mathrm{yellow}$ and $c(x_{2})=\mathrm{red}$, as depicted in Figure 12.
+Finally, we would like to introduce the last node $x_{3}$ with the hyperedges $e_{1}^{3}=\{j_{1}^{5},j_{1}^{6},x_{3}\}$, $e_{2}^{3}=\{j_{2}^{5},j_{2}^{6},x_{3}\}$ and $f=\{x_{1},x_{2},x_{3}\}$.
+For each online color, there exists a hyperedge $e\in\{e_{1}^{3},e_{2}^{3},f\}$ with $x_{3}\in e$ that already attains a makespan of $2$.
+Therefore, the node $x_{3}$ cannot be placed without increasing the makespan to $3$.
+On the other hand, extending the coloring $c$ by $c(v_{n})=\mathrm{blue}$, as suggested by Figure 12, yields a feasible coloring of the nodes, suggesting that the offline partitioning given by the colors attains a makespan of $1$ in every scenario.
+As a consequence, we obtain:
 Lemma 4.4.
@@
 […]
-Figure 4: All fifteen relevant nodes together with the hyperedges within them.
-The heights of the vertices encode their online color ($J_{1}$, $J_{2}$, $J_{3}$), while their printed colors denote their offline color.
-Any assignment of the node $x_{3}$, given the other nodes and induced hyperedges, leads to a makespan of $3$.
-Subhypergraph $\mathcal{S}$ on Three Online Colors: The Palettes.
+Figure 12: All fifteen relevant nodes together with the hyperedges within them. Notice that any assignment of the node $v_{n}$ leads to a makespan of $3$.
+4.3 Subgraph $\mathcal{S}$ on Three Online Colors: The Palettes
 It remains to consider the case where copies of the subhypergraph $\mathcal{S}$ are partitioned among the three colors, i.e., its nodes $\{1,\ldots,7\}$ are partitioned in nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$.
 Lemma 4.5.
-For every online coloring $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$ of the nodes $V(\mathcal{S})=\{v_{1},\ldots,v_{7}\}$ for nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$, there is an offline coloring $c\colon V(\mathcal{S})\to\{\mathrm{red},\mathrm{blue},\mathrm{yellow}\}$ such that two vertices exist with different online colors but the same offline color.
+For every online coloring $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$ of the nodes $V(\mathcal{S})=\{v_{1},\ldots,v_{7}\}$ for nonempty sets $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$, there is an offline coloring $c\colon V(\mathcal{S})\to\{\mathrm{red},\mathrm{blue},\mathrm{yellow}\}$ such that two nodes exist with different offline colors but the same online color.
+Proof 4.6.
+We fix the offline coloring in Figure 10, i.e., nodes $3,4$ are red, nodes $2,7$ are blue and nodes $1,5,6$ are yellow.
+The only case where the claim does not hold for the coloring is when each $J_{i}$ consists of nodes of precisely one offline color.
+In other words, up to a permutation of the colors, we have $J_{1}=c^{-1}(\{\mathrm{red}\})$, $J_{2}=c^{-1}(\{\mathrm{blue}\})$, $J_{3}=c^{-1}(\{\mathrm{yellow}\})$.
+In this case, we alter the offline coloring as follows: $1$ and $5$ are blue instead of yellow, and $2$ is yellow instead of blue.
+It is straightforward to check that this is a feasible coloring, see Figure 13.
+33112255446677
+Figure 13: The second case in the proof of Lemma 4.5. Names of hyperedges are left out for brevity.
 In our construction, we exploit this property as follows: Let us consider a copy $\mathcal{S}(t)$ and fix two of its nodes $v_{i}(t)\in J_{i}$, $v_{i^{\prime}}(t)\in J_{i^{\prime}}$ provided by Lemma 4.5 as well as a third node $v_{i^{\prime\prime}}(t)\in J_{i^{\prime\prime}}$, where $i^{\prime\prime}\notin\{i,i^{\prime}\}$.
@@
 In this case, the subhypergraph $\mathcal{S}(t)$ is called a $C$ palette (i.e., yellow palette, blue palette or red palette) and the corresponding nodes $v_{i}(t)$ are called palette nodes.
-In particular, $c(v_{i}(t))\neq C$ holds for the palette nodes of a $C$ palette.
+In particular, $c(v_{i}(t))\neq C$.
 The choice of the name is motivated by the fact that, if we connect a new node $w$ to e.g.
 yellow palette nodes, we are still allowed to extend the offline coloring by $c(w)=\mathrm{yellow}$, while maintaining that the makespan of the offline coloring is $1$.
-The Construction.
+4.4 The Construction
+Below is a summary of our construction for the case where Lemma 4.4 is not applicable.
+Essentially, we create copies of the subgraph $\mathcal{S}$ which we consecutively connect with the existing hypergraph in varying ways.
 Below is a summary of our construction for the case where Lemma 4.4 is not applicable.
@@
 For this $s$, we define $v_{\mathcal{B}}\coloneqq v_{i}(s+1)$ and $w_{\mathcal{B}}\coloneqq w_{\mathcal{B}s}$.
-[…]
-    Figure 5: An example constellation of the three palettes together with the nodes $w_{\mathcal{B}t}$.
-The $y$-axis in the palettes denote the online colorings of the palette nodes respectively.
-The offline colors of the palette nodes might look differently, but they are never the same color as the palette, which is depicted as the color of the rectangles.
-The online coloring of the nodes $w_{\mathcal{B}i}$ are not yet known.
   5. (v)
@@
 […]
+Figure 14: The introduction of the last node $v_{n}$ leads to a makespan of $3$.
+Only the seven relevant nodes are shown, although the entire hypergraph has both an online coloring and a feasible offline coloring.
+Step (i) is straightforward.
+To achieve the desired colors of $\mathcal{S}(1),\ldots,\mathcal{S}(7)$ as in step (ii), we find out the color $C$ of each $\mathcal{S}(t)$ (as a palette), and postcompose the coloring $c|_{V(\mathcal{S}(t))}\colon V(\mathcal{S}(t))\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ with the transposition
+$\tau\colon\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$
+that swaps the color $C$ with the designated color of $\mathcal{S}(t)$ as stated in Step (ii).
+Step (iii): the subinstance $\mathcal{A}$.
+Here, it suffices to connect the yellow palette $\mathcal{S}(1)$ to a new node $w_{\mathcal{A}}$ with the hyperedges $e=\{w_{1},v_{i}(1)\}$ for $t\in\{1,2,3\}$.
+By the definition of the palette nodes $v_{i}(1)$, they are not yellow.
+Since the new node $w_{\mathcal{A}}$ is only connected to non-yellow nodes, we can extend the offline coloring by $c(w_{\mathcal{A}})=\mathrm{yellow}$, maintaining its feasibility.
+For the $i_{\mathcal{A}}$ which holds $w_{\mathcal{A}}\in J_{i_{\mathcal{A}}}$, we have $e=\{w_{\mathcal{A}},v_{i_{\mathcal{A}}}(1)\}\in E(G[J_{i}])$ in particular.
+Step (iv): the subinstance $\mathcal{B}$.
+In this stage of the construction, we use the three palettes $\mathcal{S}(2)$, $\mathcal{S}(3)$, $\mathcal{S}(4)$ with palette nodes $v_{i}(t)$ for $i\in\{1,2,3\}$ for each copy $t\in\{2,3,4\}$.
+Note that we are utilizing precisely one palette of each color.
+Now we introduce three new nodes $w_{\mathcal{B}1}$, $w_{\mathcal{B}2}$ and $w_{\mathcal{B}3}$.
+These nodes are to be understood as candidates to the node $w_{\mathrm{B}}$ that we would like to define.
+They are connected to each other as well as to the three $v_{i}(t)$ of their respective palette, i.e., there exists the hyperedge $\{w_{\mathcal{B}1},w_{\mathcal{B}2},w_{\mathcal{B}3}\}\in E$ as well as hyperedges $\{w_{\mathcal{B}t},v_{i}(t+1)\}$ for $t\in\{1,2,3\}$ and $i\in\{1,2,3\}$.
+Accordingly, we may tentatively color the nodes as $c(w_{\mathcal{B}1})=\mathrm{yellow}$, $c(w_{\mathcal{B}2})=\mathrm{blue}$ and $c(w_{\mathcal{B}3})=\mathrm{red}$.
+See Figure 15 for a sketch of the introduced subhypergraph.
+$v_{2}(2)$$v_{1}(2)$$v_{2}(3)$$v_{1}(3)$$v_{3}(3)$$v_{3}(2)$$v_{1}(4)$$v_{2}(4)$$v_{3}(4)$$\mathcal{S}(2)$$\mathcal{S}(3)$$\mathcal{S}(4)$$w_{\mathcal{B}1}$$w_{\mathcal{B}2}$$w_{\mathcal{B}3}$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$
+Figure 15: An example constellation of the three palettes together with the nodes $w_{\mathcal{B}t}$.
+The $y$-axis in the palettes denote the online colorings of the palette nodes respectively.
+The offline colors of the palette nodes might look differently, but they are never the same color as the palette, which is depicted as the color of the rectangles.
+The online coloring of the nodes $w_{\mathcal{B}i}$ are not yet known.
+Now we would like to choose the node $w_{\mathcal{B}}$ ultimately.
+To this end, we consider any online assignment of the three candidates $w_{\mathcal{B}t}$ ($t=1,2,3$).
+If $\varphi(w_{\mathcal{B}1})=\varphi(w_{\mathcal{B}2})=\varphi(w_{\mathcal{B}3})$, we obtain a makespan of $3$ and can immediately truncate the instance, as it no longer has a competitive ratio better than $3$.
+Hence, there are at least two online colors $i\in\{1,2,3\}$ that the candidate nodes $w_{\mathcal{B}t}$ are assigned to, and at least one of them is not $i_{A}$.
+We pick such an online color and call it $i_{B}$.
+Step (v): the subinstance $\mathcal{C}$.
+To ensure that an edge $e_{i^{\prime\prime}}\in E(G[J_{i^{\prime\prime}}])$ is created in precisely the online color $i^{\prime\prime}\notin\{i,i^{\prime}\}$, we introduce a number of candidate nodes $w_{\mathcal{C}s}$ similar to the previous step.
+Here, it is integral that the nodes $w_{\mathcal{C}s}$ are connected to each other in a way that is
+  1. (i)
+    dense enough so that they are forced to utilize all three online colors, but also
+  2. (ii)
+    sparse enough that the offline coloring can be extended to the $w_{\mathcal{C}s}$.
+Conveniently, we can recollect to our considerations about the subgraph $\mathcal{S}$: If we create seven copies of it, one of them is forced to attain all three online colors, or else we can complete the hypergraph in a way that no online coloring is better than $3$-competitive.
+Recalling that we may connect a node $w_{j}$ with the three nodes $v_{i}(t)$ of the palette in its offline color, we define the seven nodes as depicted in Figure 10.
+More precisely, we create palettes $\mathcal{S}(5)$, $\mathcal{S}(6)$, $\mathcal{S}(7)$ and introduce nodes $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$ together with the following hyperedges:
+  • $\{w_{\mathcal{C}1},w_{\mathcal{C}2},w_{\mathcal{C}3}\}$, $\{w_{\mathcal{C}2},w_{\mathcal{C}4},w_{\mathcal{C}5}\}$, $\{w_{\mathcal{C}4},w_{\mathcal{C}6},w_{\mathcal{C}7}\}$, $\{w_{\mathcal{C}3},w_{\mathcal{C}6}\}$, $\{w_{\mathcal{C}3},w_{\mathcal{C}7}\}$, as well as their subsets of size $2$ (so that $G[\{w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}\}]\cong\mathcal{S}$),
+  • $\{v_{i}(5),w_{\mathcal{C}s}\}$ for $i\in\{1,2,3\}$, $s\in\{1,5,6\}$, i.e., between yellow palette nodes and new nodes colored yellow in Figure 10,
+  • $\{v_{i}(6),w_{\mathcal{C}s}\}$ for $i\in\{1,2,3\}$, $s\in\{2,7\}$, i.e., between blue palette nodes and new nodes colored blue in Figure 10,
+  • $\{v_{i}(7),w_{\mathcal{C}s}\}$ for $i\in\{1,2,3\}$, $s\in\{3,4\}$, i.e., between red palette nodes and new nodes colored red in Figure 10.
+An overview of the construction, without specifying the online colors of the nodes $w_{\mathcal{C}s}$, is given in Figure 16.
+$v_{1}(1)$$v_{2}(1)$$v_{3}(1)$$v_{1}(2)$$v_{2}(2)$$v_{3}(2)$$v_{1}(3)$$v_{2}(3)$$v_{3}(3)$$w_{\mathcal{C}3}$$w_{\mathcal{C}1}$$w_{\mathcal{C}2}$$w_{\mathcal{C}5}$$w_{\mathcal{C}4}$$w_{\mathcal{C}6}$$w_{\mathcal{C}7}$$\mathcal{S}(5)$$\mathcal{S}(7)$$\mathcal{S}(6)$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$$J_{1}$$J_{2}$$J_{3}$
 Figure 16: Partial sketch of subinstance $\mathcal{C}$.
-Notice that we have maintained an offline coloring of the nodes with makespan $1$ so far up to the very last node, and the last node $n$ neighbors only blue and yellow nodes by construction and therefore the offline assignment $c(n)=\mathrm{red}$ does not increase the makespan.
-In total, this yields that the offline optimal value is $1$.
-A careful enumeration of introduced hyperedges reveals that $233$ hyperedges suffice to create a monochromatic hyperedge for any of the cases demonstrated, which proves Theorem 4.1.
-A Note on Theorem 1.2.
-Here, we abstain from presenting a proof of Theorem 1.2.
-Our construction in this general case is also based on incrementally creating monochromatic edges of each online color.
-However, we are faced with two fundamental challenges:
+Recall that our goal is to find an edge of the form $\{v_{i^{\prime\prime}}(t),w_{\mathcal{C}s}\}\subseteq J_{i^{\prime\prime}}$ for the $i^{\prime\prime}\notin\{i,i^{\prime}\}$.
+If the nodes $w_{\mathcal{C}s}$ for $s\in\{1,\ldots,7\}$ are spread across the three online colors, there is an index $s\in\{1,\ldots,7\}$ with the property that $w_{\mathcal{C}s}\in J_{i^{\prime\prime}}$; for a suitable $t$ depending on the offline color of the node $w_{\mathcal{C}s}$, we thus find the node $v_{i^{\prime\prime}}(t)$ and hence the aforementioned edge as desired.
+In order to guarantee the existence of $w_{\mathcal{C}s}\in J_{i^{\prime\prime}}$, we need to convince ourselves that the subgraph induced by $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$ is partitioned on all three online colors.
+Conveniently, we have seen in Subsection 4.2 that if we create seven copies of the subgraph $G[\{w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}\}]$, one of the copies must admit all three online colors or else the assignment cannot be better than $3$-competitive by Lemma 4.4.
+Therefore, we create seven copies of the subgraph induced by $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$, combining the nodes in each copy with the corresponding palette nodes.
+In total, one of the seven copies of $w_{\mathcal{C}1},\ldots,w_{\mathcal{C}7}$ must admit all three online colors, in particular the online color $i^{\prime\prime}$.
+For a node $w_{\mathcal{C}s}$ attaining this online color, we find a palette node $v_{i^{\prime\prime}}(t)$ such that $\{v_{i^{\prime\prime}}(t),w_{\mathcal{C}s}\}\in J_{i^{\prime\prime}}$.
+Now we compose the offline coloring $c$ obtained so far with a permutation of the three offline colors such that $c(v_{i^{\prime\prime}}(t))=\mathrm{blue}$ and $c(w_{\mathcal{C}s})=\mathrm{yellow}$.
+The reader might wonder why we need all three components; indeed, we will have three hyperedges of the type $e_{i}$ already in the underlying hypergraph of $\mathcal{C}$ alone.
+However, we require that the hyperedges $e_{i}$ are each colored in blue and yellow endpoints, which is not possible when we take all of them from $\mathcal{C}$.
+Step (vi): Recoloring the connected components.
+Finally, we permute the colors in the three connected components created in Steps (iii), (iv) and (v) (as shown in Figure 15) such that $c(w_{\mathcal{B}{i_{\mathcal{B}}}})=\mathrm{yellow}$ and $c(v_{i_{\mathcal{B}}}(i_{\mathcal{B}}))=\mathrm{blue}$.
+See Figure 17 for an example of the permutation process.
+$v_{1}(4)$$v_{2}(4)$$v_{3}(4)$$\mathcal{S}(4)$$w_{\mathcal{B}3}$112233
+Figure 17: Permuting the colors of the palette $\mathcal{S}(4)$.
+This permutation was implied by the assignment $\varphi(w_{\mathcal{B}3})=i_{\mathcal{B}}=3$.
+We have to ensure that $c(v_{3}(3))=\mathrm{blue}$ and $c(w_{\mathcal{B}3})=\mathrm{yellow}$, which is achieved by transposing the offline colors yellow and blue.
+Of course, this transposition is applied to the entire component, such that e.g.
+$c(w_{\mathcal{B}1})=\mathrm{blue}$, which is omitted here.
+We are now ready to execute Step (vii) and force instance to admit a makespan of $3$. To this end, we summarize our setting so far.
+Lemma 4.7.
+There exists a family $\mathcal{I}$ of instances over the node set $V$ together with feasible colorings $c\colon V\to\{\mathrm{blue},\mathrm{red},\mathrm{yellow}\}$ such that for every online assignment $\varphi\colon V\to\{1,2,3\}$, at least one instance $I\in\mathcal{I}$ satisfies one of the following properties:
   1. (i)
-    The monochromatic hyperedges increase in size rather slowly, yet unless they have size of $m-1$, they are not strong enough to act as a palette for other hyperedges in the same sense as here.
+    $\mathrm{makespan}(I)=3$, or
   2. (ii)
-    To use our construction for Corollary 1, we restrict to hypertrees. This is a significant constraint, as it roughly implies that every palette may be used at most once.
-Therefore, we include a more intricate recursive construction with much larger hypergraphs in the full version of the paper.
+    there exist neighboring nodes $v_{i}(t),w_{i}\in J_{i}$ for each $i\in\{1,2,3\}$ such that $c(v_{i}(t))=\mathrm{blue}$ and $c(w_{i})=\mathrm{yellow}$.
+Step (vii): the last red node
+As sketched in the introduction of this subsection (cf.
+Figure 14 in particular), we insert the very last node $n$ by combining it with the nodes $v_{i}(t)$ and $w_{i}$.
+More precisely, this last node is incident to the hyperedges $\{v_{i}(t),w_{i},n\}$ for $i\in\{1,2,3\}$.
+No matter which online color it is assigned to, a monochromatic hyperedge of that particular online color emerges.
+Notice that we have maintained a feasible offline coloring of the nodes so far up to the very last node, and the last node $n$ neighbors only blue and yellow nodes by construction and therefore the offline assignment $c(n)=\mathrm{red}$ extends the feasible coloring.
+In total, this yields that the offline optimal value is $1$, implying Theorem 4.1.
+A careful enumeration of introduced hyperedges reveals that $233$ hyperedges suffice for any of the cases demonstrated, which we present next.
+4.5 Analysis
+The motivation for the aforementioned family of instances is that, already for a finite number $K$ of hyperedges, the competitive ratio is no better than the trivial bound of $3$.
+Although it is not necessarily the case that our construction has the minimal possible size, we would still like to analyze it so as to provide an upper bound to the smallest number of scenarios for which no algorithm with a non-trivial competitive ratio exists.
+To recollect our construction, we create seven copies of the subgraph $\mathcal{S}$ and apply a case distinction:
+  1. (A)
+    If every copy is partitioned among at most two online colors as in Subsection 4.2, we insert three additional nodes $x_{1}$, $x_{2}$, $x_{3}$ and seven new hyperedges to obtain an adversarial instance.
+Together with the seven copies of $\mathcal{S}$, which have $7$ nodes and $14$ hyperedges each, this instance adds up to $7\cdot 7+3=52$ nodes and $7\cdot 14+7=105$ hyperedges.
+  2. (B)
+    If the above case does not hold, the seven steps given previously are executed in order to create an adversarial instance. To be more precise:
+    • Throughout the execution of the seven steps, at most $6$ of the copies are assumed to be partitioned onto exactly two online colors, which means up to $6\cdot 7=42$ nodes and $6\cdot 14=84$ hyperedges.
+    • In addition to those copies, there are $7$ copies that are needed for Step (i), i.e., $7\cdot 7=49$ nodes and $7\cdot 14=98$ hyperedges.
+    • In Step (iii), we introduce one more node $w_{\mathcal{A}}$ with $3$ new hyperedges.
+    • In Step (iv), we introduce 33 new nodes $w_{\mathcal{B}i}$ $(i\in\{1,2,3\})$.
+Each new node is connected to a palette by 33 new hyperedges of size $2$ and to each other with one hyperedge of size $3$, meaning that Step (iv) adds $3\cdot 3+1=10$ new hyperedges.
+    • Step (v) essentially adds an eighth copy of $\mathcal{S}$, i.e., $7$ nodes and $14$ hyperedges, and connects this copy to the existing instance using three hyperedges of size $2$ per inserted node, which adds up to $7\cdot 3=21$ more hyperedges; i.e., $14+21=35$ for the entire step.
+    • Finally, there is the last node that leads to the competitive ratio of $3$, it is connected to the existing instance by $3$ hyperedges.
+In total, we have introduced up to $42+49+1+3+7+1=103$ nodes and up to $84+98+3+10+35+3=233$ hyperedges, as claimed in Theorem 4.1.
+5 Tight Bounds for Arbitrary $m$ and Sufficiently Many Scenarios
+Our main goal in this section is to prove the following theorem.
+Theorem 5.1.
+Let $m\in\mathbb{N}_{\geq 1}$.
+There exists a number $N(m)$ such that there is no $(m-\varepsilon)$-competitive algorithm for any $\varepsilon>0$ for Online Makespan Hypergraph Coloring, even when restricted to hyperforests with $N(m)\leq(2m+1)\uparrow\uparrow 5$ (i.e., a tower of height $5$) nodes of dimension at most $m$.
+In terms of scheduling, this translates to the following statement.
+Corollary 5.2.
+Let $m\in\mathbb{N}_{\geq 1}$.
+There exists a number $K_{m}$ such that there is no $(m-\varepsilon)$-competitive algorithm for any $\varepsilon>0$, for Online Makespan Scheduling under $K$ Scenarios for $K\geq K_{m}$, even when restricted to unit processing times and $|S_{k}|\leq m$ for every scenario $S_{k}\in\{S_{1},\ldots,S_{K_{m}}\}$.
+One might wonder whether the number $K_{m}$ can be strengthened to a global $K$ for all possible numbers $m$ of machines.
+However, the generalization of Graham’s List Scheduling Algorithm is $\max\{m,K+1\}$-competitive, even on varying processing times, answering this question negatively.
+The reader might further notice the subtlety that Theorem 1.2 and Corollary 5.2 are given in terms scenarios, while Corollary 1 and Theorem 5.1 are given in terms of nodes (which normally correspond to jobs).
+This is a deliberate choice; indeed, while our problem on scheduling parametrizes the makespan over the number $K$ of scenarios instead of the number $n$ of jobs, it is more common in hypergraph coloring literature to express the number $m$ of colors in terms of the number $n$ of nodes instead of the number $K$ of hyperedges.
+Moreover, as it is easier to count nodes in our construction, we do so throughout this section and converse our bounds to the number of hyperedges at the very end.
+In an attempt to create a badly-behaving family of instances, let us turn the problem around: Now, we become the adversary and reveal for every new node $j\in[n]$ some hyperedges $e\subseteq[j]$ with the following properties:
+  1. (i)
+    $j\in e$,
+  2. (ii)
+    $e\setminus\{j\}$ has been revealed as a hyperedge before, unless $e=\{j\}$.
+The algorithm then assigns the new node to some online color, which we have no control over.
+As the adversary, we are once again interested in maintaining a feasible offline coloring in addition to the online coloring determined by the algorithm.
+Since we define our instance recursively, this is a much more complicated task.
+Therefore, we would like to maintain an active subhypergraph $\mathcal{G}$ of our hypergraph $\mathcal{H}$ such that every connected component of $\mathcal{G}$ consists of exactly one maximal hyperedge, and these hyperedges are pairwise disconnected in the hypergraph $\mathcal{H}$.
+Accordingly, we also call the hyperedges $e\in E(\mathcal{G})$ and their nodes $v\in e$ active, as well, as long as they are part of the active subhypergraph.
+Active hypergraphs serve a crucial purpose in our construction.
+If we reveal a new node $j\in[n]$ incident to multiple maximal hyperedges, at most one of these hyperedges can remain active, since they are no longer disconnected in $\mathcal{H}$.
+Throughout our construction, we always connect new nodes the existing hypergraph via extending active hyperedges.
+This way, we ensure that the entire hypergraph $\mathcal{H}$ remains a hyperforest, hence a feasible offline coloring with $d\leq m$ colors exists where $d$ is the maximum number of nodes of a hyperedge.
+Inspired by the case distinction of hyperedges being split among two or three online colors in the previous construction, we describe a milestone which we intend to repeat to force larger monochromatic subgraphs.
+Lemma 5.3.
+For Online Makespan Hypergraph Coloring(m)(m), there exists a family of instances $\mathcal{I}(m,d)$ with at most $X(m,d)$ nodes for a suitable number $X(m,d)$ each, on which the adversary achieves against any algorithm an online coloring $\varphi\colon[n]\to[m]$ with at least one of the following properties:
+  1. (i)
+    For every online color $i\in[m]$, there is a node of color $i$, i.e., it holds that $\varphi^{-1}(\{i\})\neq\emptyset$. Moreover, there is at least one node of each online color that is active.
+  2. (ii)
+    For an online color $i\in[m]$, there exists an active monochromatic hyperedge of size $d$, i.e., $e\in E(\mathcal{H})$, $|e|=d$ and $\varphi(j)=i$ for every $j\in e$.
+The number $X(m,d)$ is implicitly defined by the recursion
+$X(m,d)=\begin{cases}1,&d=1\text{ or }m=1\\ X(m-1,d)\cdot((m-1)\cdot X(m,d-1)+1)+X(m,d-1),&\text{else}\end{cases}$
+Proof 5.4.
+We prove this statement by induction. Evidently for $m=1$, $X(1,d)=1$ node suffices to achieve the first property. Similarly for $d=1$, $X(m,1)=1$ node suffices to achieve the second property.
+Now, we fix $(m,d)$ and assume that the statement holds for all $1\leq m^{\prime}\leq m$, $1\leq d^{\prime}\leq d$, $(m^{\prime},d^{\prime})\neq(m,d)$.
+First, we observe by induction that we know an instance $\mathcal{I}(m,d-1)$ with X(m,d1)X(m,d-1) nodes that either has a node of every online color, or a monochromatic (d1)(d-1)-dimensional hyperedge.
+If the former property holds, then the statement follows for $(m,d)$ as well.
+Therefore, we may assume without loss of generality that, the second property holds every time we reveal a copy of this instance (in the same order of nodes as revealed in the inductive assumption).
+We execute $X(m-1,d)\cdot(m-1)+1$ copies of the instance $\mathcal{I}(m,d-1)$.
+Here, the number X(m1,d)X(m-1,d) is finite due to the inductive assumption.
+From each instance, we maintain solely the (d1)(d-1)-dimensional hyperedge (which exists by the second property of the lemma) as active.
+Each such hyperedge is assigned an online color.
+If there exists a hyperedge of every online color, then the first property holds and the statement is shown.
+Therefore, there are hyperedges of at most (m1)(m-1) colors.
+By the pigeonhole principle, there is an online color (the online color $m$, without loss of generality) such that $\varphi^{-1}(\{m\})$ contains at least X(m1,d)+1X(m-1,d)+1 pairwise node-disjoint hyperedges of size d1d-1.
+We have now secured X(m1,d)+1X(m-1,d)+1 active hyperedges $e_{1},\ldots,e_{X(m-1,d)+1}$ which we use to simulate OMHC(m1)(m-1).
+To this end, we reveal a copy of the instance $\mathcal{I}(m-1,d)$.
+In addition to the hyperedges within $\mathcal{I}(m-1,d)$, we reveal for the $j$-th node the hyperedge $e_{j}\cup\{j\}$.
+If $j$ is assigned the online color $m$, the second property holds and our proof is completed.
+Therefore, we may assume that all nodes of the copy of $\mathcal{I}(m-1,d)$ are assigned one of the online colors $\{1,\ldots,m-1\}$.
+After every such assignment, exactly one hyperedge $e_{j}$ is deactivated, so that the (d1)(d-1)-dimensional hyperedge $e_{X(m-1,d)+1}$, which is monochromatic of color $m$, remains active.
+Now we observe: Since the subhypergraph induced by the last X(m1,d)X(m-1,d) nodes is isomorphic to $\mathcal{I}(m-1,d)$ and the nodes must be partitioned among m1m-1 colors, one of the two properties hold; either there is a monochromatic $d$-dimensional hyperedge, or all online colors $1,\ldots,m-1$ are attained (as is $m$, due to the hyperedges).
+In either of the cases, the property follows for the entire instance, which we define as $\mathcal{I}(m,d)$.
+We create at most $X(m-1,d)\cdot(m-1)+1$ copies of at most X(m,d1)X(m,d-1) nodes each, followed by at most another X(m1,d)X(m-1,d) nodes. In total, this accounts to at most
+$(X(m-1,d)\cdot(m-1)+1)\cdot X(m,d-1)+X(m-1,d)$
+$=$ $X(m-1,d)\cdot((m-1)\cdot X(m,d-1)+1)+X(m,d-1)$
+nodes, as claimed.
+Finding a closed form of the above recursion turns out to be challenging, though we provide a rather rough estimate for the curious reader’s intuition as well as for later use.
+Lemma 5.5.
+For $X(m,d)$ as in Lemma 5.3, we have $X(m,d)\leq\prod_{i=1}^{m+d}(m+d)^{2^{i}}$.
+Proof 5.6.
+We prove this by induction. For $m=1$ and $d=1$, the claim is vacuously true. Assuming that it holds for $1\leq d^{\prime}\leq d$, $(m^{\prime},d^{\prime})\neq(m,d)$, we compute
+$X(m,d)$ $=X(m-1,d)\cdot((m-1)\cdot X(m,d-1)+1)+X(m,d-1)$
+$\leq(m+1)\cdot X(m-1,d)\cdot X(m,d-1)$
+$\leq(m+1)\cdot\prod_{i=1}^{m+d-1}(m+d)^{2^{i}}\cdot\prod_{i=1}^{m+d-1}(m+d)^{2^{i}}$
+$\leq(m+1)\cdot\prod_{i=1}^{m+d+1-i}\left((m+d)^{2^{i}}\right)^{2}$
+$=(m+1)\cdot\prod_{i=1}^{m+d-1}(m+d)^{2^{i+1}}$
+$\leq(m+1)\cdot\prod_{i=1}^{m+d}(m+d)^{2^{i}}$
+We can further bound in Lemma 5.5 as
+$X(m,d)\leq(m+1)\cdot\prod_{i=1}^{m+d}(m+d)^{2^{i}}\leq(m+1)(m+d)^{2^{m+d+1}}.$
+In particular,
+$X(m,m)\leq(m+1)\cdot(2m)^{2^{2m+1}}.$ (30)
+Both of the properties we force to exist in Lemma 5.3 prove useful for our adversarial purposes.
+In particular, if we let $m=d$, the second property implies that the algorithm is no better than $m$-competitive, which we want to show.
+However, we cannot exclude the first property, which we also want to lead to an incompetitive demise in the sequel.
+Having been granted a hypergraph $\mathcal{H}$ with a monochromatic active node of each online color, we proceed similarly as in the case of $m=3$, in that the many copies of $\mathcal{H}$ act as “palettes”.
+However, now we must also grow the palettes, i.e., the size of the palettes must increase throughout our construction.
+This is especially challenging because we can only extend active hyperedges (or in general, at most one hyperedge per connected component).
+We try to increase monochromatic hyperedges in size by revealing sufficiently many copies of $\mathcal{I}(m,m)$.
+Lemma 5.7.
+For OMHC(m)(m), there exists a family of instances $\mathcal{L}(m,d)$ which are hyperforests with at most $Y(m,d)$ nodes each for a suitable number $Y(m,d)$, such that the adversary achieves on $\mathcal{L}(m,d)$ at least one of the following properties against any algorithm resulting in an online coloring $\varphi\colon[n]\to[m]$:
+  1. (i)
+    For every online color $i\in[m]$, there are $d$-dimensional hyperedges $e_{1},\ldots,e_{m}$ whose nodes are all active and the nodes in $e_{i}$ have the online color $i$.
+Furthermore, these hyperedges are pairwise disconnected in $\mathcal{L}(m,d)$.
+  2. (ii)
+    For an online color $i\in[m]$, there exists an active monochromatic hyperedge of size $m$, i.e., $e\in E(\mathcal{L}(m,d))$, $|e|=m$ and $\varphi(j)=i$ for every $j\in e$.
+The number $Y(m,d)$ is implicitly defined by the recursion
+$\hskip-8.5359ptY(m,d)=\begin{cases}1,&m=1\\ X(m,m)&d=1\\ m^{2}\cdot(Y(m-1,d)^{m}+1)\cdot(Y(m,d-1)+1)+m\cdot Y(m-1,d)^{m+1}+1,&\text{else}.\end{cases}$
+Proof 5.8.
+As in the previous proof, we prove the statement by induction.
+For $d=1$, we can set $\mathcal{L}(m,1)=\mathcal{I}(m,m)$ and $Y(m,1)=X(m,m)$.
+For $m=1$, $Y(1,d)=1$ nodes suffices to satisfy the second property.
+Let us assume that the supposed instances $\mathcal{L}(m^{\prime},d^{\prime})$ with $Y(m^{\prime},d^{\prime})$ nodes each exist for every $m^{\prime}\leq m,d^{\prime}\leq d,(m^{\prime},d^{\prime})\neq(m,d)$.
+We first release
+$a_{1}\coloneqq m^{2}\cdot Y(m-1,d)^{m}+1$ (31)
+disjoint copies of $\mathcal{L}(m,d-1)$.
+If any of the copies satisfy the second property, the statement follows immediately.
+Therefore, we may assume that each copy satisfies the first property, i.e., we obtain $a_{1}$ hyperedges of each color.
+By virtue of the first property, the hyperedges within a copy are pairwise disconnected, so that we can mark all of them as active.
+Next, we release $a_{1}$ singleton nodes.
+These extend $m$ hyperedges per node; one in each color.
+Therefore, they extend precisely one monochromatic hyperedge of size d1d-1 to a monochromatic hyperedge of size $d$ per node.
+We mark the new monochromatic hyperedges as active and all others as inactive.
+Unfortunately, the online algorithm assigns the singleton nodes, which we have no control over.
+However, we can guarantee that in one of the colors, in the $m$-th color without loss of generality, there are at least
+$a_{2}\coloneqq m\cdot Y(m-1,d)^{m}+1$ (32)
+active hyperedges $e_{1},\ldots,e_{a_{2}}$.
+Then, we release $m\cdot Y(m-1,d)^{m-1}$ copies of $\mathcal{L}(m-1,d)$.
+We connect these copies to our existing active hyperedges as follows: In addition to the hyperedges within $\mathcal{L}(m-1,d)$, each node in the copy extends one of the hyperedges $e_{k}$, e.g., one with the smallest size that is active, excluding $e_{a_{2}}$.
+This way, every node extends one active hyperedge of size (d1)(d-1) of colors $1,\ldots,m-1$ each.
+Let us call a copy of $\mathcal{L}(m-1,d)$ successful if the algorithm assigns no node $v$ in this copy the color $m$, and unsuccessful otherwise.
+Claim 6.
+There can be at most $m\cdot Y(m-1,d)^{m}-1$ unsuccessful copies unless the second property is satisfied.
+Proof 5.9 (Proof of Claim 6).
+For the sake of contradiction, let us assume that there are $m\cdot Y(m-1,d)^{m}$ unsuccessful copies.
+Consider the first $m\cdot Y(m-1,d)^{m}$ of these copies.
+For every unsuccessful copy, at least one active hyperedge of size $d$ must be extended to size $d+1$ and hence may stay active, and in return, Y(m1,d)1Y(m-1,d)-1 hyperedges must become inactive.
+Since we always pick the smallest-size hyperedges to extend, the first $Y(m-1,d)^{m-1}$ copies exhaust all hyperedges $e_{k}$ (except $e_{a_{2}}$ which we ignore until the very end of our construction).
+In the process, $Y(m-1,d)^{m-1}$ hyperedges stay active, which are used once again for the next $m\cdot Y(m-1,d)^{m}$ copies and so on.
+After the $d^{\prime}$-th run of this procedure for $d^{\prime}\leq m-d$, we have $Y(m-1,d)^{d+m-d^{\prime}}$ active hyperedges of size $d+d^{\prime}$.
+At the end, for $d^{\prime}=m-d$, we have one hyperedge of size $m$, satisfying the second property.
+By the claim above, at least one copy must be successful, i.e., partitioned among colors $1,\ldots,m-1$.
+Once we find a successful copy, the inductive assumption on $\mathcal{L}(m-1,d)$ delivers one $d$-dimensional hyperedge of colors $1,\ldots,m-1$ each, which are pairwise disconnected.
+Together with $e_{a_{2}}$, this yields the first property overall.
+It remains to count the numbers of nodes and hyperedges.
+  1. (i)
+    We first added $a_{1}$ copies of $\mathcal{L}(m,d-1)$, which accounts to $a_{1}\cdot Y(m,d-1)$ nodes.
+  2. (ii)
+    Then, we added $a_{1}$ singleton nodes.
+  3. (iii)
+    Lastly, we released $a_{2}-1$ copies of $\mathcal{L}(m-1,d)$ together with a very last node, which accounts to $a_{2}\cdot Y(m-1,d)+1$ nodes.
+In total, we utilized
+$a_{1}\cdot(Y(m,d-1)+1)+a_{2}\cdot Y(m-1,d)+1$ (33)
+$=$ $m^{2}\cdot(Y(m-1,d)^{m}+1)\cdot(Y(m,d-1)+1)+m\cdot Y(m-1,d)^{m+1}+1$ (34)
+nodes, as claimed.
+It is worth noting that the terms in the above proof are not tight: For the sake of simplicity, we have rounded up some terms without changing the asymptotics of the result.
+The previous lemma almost immediately implies Theorem 5.1, which we ultimately wanted to prove.
+Proof 5.10 (Proof of Theorem 5.1).
+We choose $N(m)\coloneqq Y(m,m-1)+1$, where $Y\colon\mathbb{N}\times\mathbb{N}\to\mathbb{N}$ is defined as in Lemma 5.7.
+Indeed, we could build the instance $\mathcal{L}(m,m-1)$, which provides pairwise disconnected hyperedges $e_{1},\ldots,e_{m}$ of each color $1,\ldots,m$ by Lemma 5.7.
+Now a last node $v$ can be introduced so that it extends $e_{1},\ldots,e_{m}$.
+For the color $i$ that the node $v$ is assigned, $e_{i}\cup\{v\}$ is a monochromatic edge of color $i$.
+By the recursion in Lemma 5.7, it follows that
+$Y(m,d)\leq(Y(m-1,d)\cdot Y(m,d-1))^{m+2}\leq\max\{Y(m-1,d)\cdot Y(m,d-1)\}^{2m+4}.$
+Extending the terms Y(m,d1)Y(m,d-1) and Y(m1,d)Y(m-1,d) and iterating this process, we obtain
+$Y(m,m-1)\leq X(m,m)^{(m+2)^{m-1}}\leq\left((m+1)\cdot{(2m)^{2^{(2m+1)}}}\right)^{(m+2)^{m-1}}$
+by (30).
+Very roughly, we can estimate $N(m)\leq(2m+1)\uparrow\uparrow 5$ (i.e., a tower of height $5$), as claimed in Theorem 5.1.
+While we are still parameterizing over nodes, let us first deduce Corollary 1.
+Theorem 5.1 reveals that we cannot color hyperforests with with $(2m+1)\uparrow\uparrow 5$ nodes in $m$ colors without creating a monochromatic hyperedge.
+In other words, if we want a feasible coloring of a hyperforest with $2n+1$ nodes, we need at least $\mathrm{slog_{5}(n)}$ colors, as claimed.
+Throughout this section, we have counted nodes instead of hyperedges as it is stated in Theorem 1.2.
+However, hyperforests with hyperedges of size at most $m$ with $n$ nodes have $O(mn)$ hyperedges, as $mn$ is the number of incidences and there cannot be more hyperedges than incidences.
+As both Theorem 1.2 and Corollary 5.2 only state the finiteness of the number $K_{m}$ of scenarios, they follow immediately from Theorem 5.1 as well.
 6 Tight Bounds For $K=3$ and Unit Processing Times
-In the previous section, we have established that for large enough $K$, there is no better-than-$3$-competitive algorithm for OMSS$(3,K)$, even for unit processing times.
-Let us reverse the parameters and ask: How competitively can we solve OMSS$(m,3)$?
-This is not an easy question, indeed, even for OMSS$(m,1)$, $m\geq 4$, a tight bound is not known.
-Restricting to the unit processing time case $(p_{j}\equiv 1)$, we again obtain a tight bound, which is the smallest possible ratio (due to Theorem 2.3):
+In the previous two sections, we have established in two ways that for large enough $K$, there is no better-than-$3$-competitive algorithm for OMSS(3,K)(3,K), even for unit processing times.
+Let us reverse the parameters and ask: How competitively can we solve OMSS(m,3)(m,3)?
+This is not an easy question, indeed, even for OMSS(m,1)(m,1), a tight bound is not known for most $m$.
+Restricting to the unit processing time case $(p_{j}=1)$, we achieve a tight bound, which is the smallest possible ratio considering the lower bound of $\rho\geq 2$ provided by Theorem 3.1.
+We present a round robin algorithm that is $2$-competitive in the special case of $p_{j}\equiv 1$ for an arbitrary number $m$ of machines.
 Theorem 6.1.
-For all $m\in\mathbb{N}$, there is a $2$-competitive algorithm for the special case of OMSS$(m,3)$ where the jobs have unit processing times.
-The algorithm in question employs an adapted round-robin method which assigns jobs to machines according to the scenarios they appear in.
-Essentially, we try to assign jobs $j\in S_{1}\setminus(S_{2}\cup S_{3})$ and $j^{\prime}\in(S_{2}\cup S_{3})\setminus S_{1}$ in a similar manner, and analogously for cyclic permutations of the indices $i\in\{1,2,3\}$.
-This gives a partition of jobs $j\in[n]\setminus(S_{1}\cap S_{2}\cap S_{3})$ in three categories.
-The round robin assignments for each category are translated by roughly $\nicefrac{{m}}{{3}}$, and jobs $j\in S_{1}\cap S_{2}\cap S_{3}$ are handled separately.
-The algorithm is best visualized by drawing an analogue to a bingo card, for which we refer the reader to the full version.
+For all $m\in\mathbb{N}$, there is a $2$-competitive algorithm for the special case of OMSS(m,3)(m,3) where the jobs have unit processing times.
+The algorithm is best visualized through an infinite sequence of “bingo cards” with $m$ rows and $3$ columns, each of which looks like Figure 18.
+12323231313121211232323131312122123232313131212312323231313121241232323131312125123232313131212612323231313121271232323131312128
+Figure 18: Bingo card for $m=8$. Although it admittedly rather resembles a tombola card turned sideways, our analysis will be better suited for the game of bingo.
+Here, every row represents a machine, the above example illustrates $m=8$.
+Each of the triangles correspond to a single job, where jobs $j\in S_{1}\cap S_{2}\cap S_{3}$ are to be understood as an entire square.
+We assign jobs only by looking at the column they belong to, where we fill the columns in a round-robin procedure.
+For the first column, the counting index of the round robin starts in the first row, for the second, it starts in the $\left(\left\lceil\frac{m}{3}\right\rceil+1\right)$-st row, and for the third, it starts in the $\left(\left\lceil\frac{2m}{3}\right\rceil+1\right)$-st.
+We count downwards until the lowest row of the column is reached, then we proceed from the top downwards.
+See Figure 19 for an example.
+[…]
+Figure 19: Counting indices of the above bingo card.
+We place each job $j\in[n]\setminus(S_{1}\cap S_{2}\cap S_{3})$ to the first triangle it fits into according to the counting index described above.
+If the job $j$ does not fit into any triangle, we create a new bingo card.
+We place a job $j\in S_{1}\cap S_{2}\cap S_{3}$ in an empty square with largest possible counting index.
+We break ties in favor of the rightmost column.
+If such a square does not exist, we create a new bingo card.
+This procedure is formalized in Algorithm 2 for the special case that all jobs fit a single bingo card.
+The general procedure is that, we find the earliest-created bingo card, and apply the lines 9-19 of Algorithm 2 on it.
+Here, the counting indices must then be maintained for every bingo card separately.
+We will first analyze instances where 2 does not assign two jobs onto the same placeholder (e.g., two jobs $j\in S_{1}\cup S_{2}\setminus S_{3}$ to the same triangle or a job $j\in S_{1}\cup S_{2}\cap S_{3}$ on top of a non-vacant triangle).
+At the end, we argue how we can reduce the general case to this special case.
+[…]
+Algorithm 2 The bingo card algorithm for $m$ arbitrary, $K=3$, $p_{j}\equiv 1$, assuming that jobs fit in a bingo card.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+8:for $j=1$ to $n$ do
+9:  Let $S\subseteq[3]$ be the set of scenario indices such that $j\in S_{k}$
+10:  if $S\neq[3]$ then
+11:   $J_{i_{S}}\leftarrow J_{i_{S}}\cup\{j\}$
+12:   $i_{S}\leftarrow i_{S}+1$
+13:   $i_{S}\leftarrow(i_{S}\mod m)$
+14:  else
+15:   $k\leftarrow\mathrm{argmax}_{k\in[3]}\{i_{\{1,2,3\}}^{k}+\nicefrac{{k}}{{4}}\}$
+16:   $J_{i_{\{1,2,3\}}^{k}}\leftarrow J_{i_{\{1,2,3\}}^{k}}\cup\{j\}$
+17:   $i_{\{1,2,3\}}^{k}\leftarrow i_{\{1,2,3\}}^{k}-1$
+18:   $i_{\{1,2,3\}}^{k}\leftarrow(i_{\{1,2,3\}}^{k}\mod m)$
+19:  end if
+20:end for
+Analysis.
+We are ready to analyze Algorithm 2 by defining our first type of bingo. Recall that the instance that we analyze consists of a single bingo card.
+Definition 7.
+A row bingo, given a solution $J_{1}\dot{\cup}J_{2}\dot{\cup}J_{3}$ is a machine $i$, for which a scenario $k\in\{1,2,3\}$ with $p(J_{i}\cap S_{k})=3$ exists.
+The following is an immediate consequence of this definition:
+Proposition 8.
+Any schedule without a row bingo is $2$-competitive.
+Proof 6.2.
+Any nonempty schedule without a row bingo has makespan at most $2$, and an offline optimum of at least $1$.
+Therefore, we may assume without loss of generality that the $m^{\prime}$-th machine is a row bingo for some $m^{\prime}=\left\lceil\frac{m_{1}\cdot m}{3}\right\rceil+m_{2}$ ($m_{1}\in\{0,1,2\}$, $m_{2}\in\{0,\ldots,\left\lceil\frac{m}{3}\right\rceil-1\}$ or $m_{1}=1,m_{2}=\left\lceil\frac{m}{3}\right\rceil,m\equiv 2\mod{3}$ or $m_{1}=2,m_{2}=\left\lceil\frac{m}{3}\right\rceil,m\equiv 0\mod{3}$).
+Our goal is to show that $p(S_{k})>m$, which implies that the offline optimum in the $k$-th scenario is at least $2$, bounding the competitive ratio by $\frac{3}{2}<2$.
+We apply a case distinction.
+To this end, we first observe zones into which a bingo card is decomposed by the round robin counting indices.
+For an example, see Figure 20.
+$m-\left\lceil\frac{2m}{3}\right\rceil$ $\left\lceil\frac{2m}{3}\right\rceil-\left\lceil\frac{m}{3}\right\rceil rows$$\left\lceil\frac{m}{3}\right\rceil$ rows12323231313121211232323131312122123232313131212312323231313121241232323131312125123232313131212612323231313121271232323131312128
+Figure 20: The same bingo card.
+The yellow colored squares were filled by at least one triangle, while the blue square has been filled by a square (i.e., a job $j\in S_{1}\cap S_{2}\cap S_{3}$).
+Notice that by the translated round-robin indexing, zones of nearly equal heights occur.
+If the first row is a row bingo for, say, the first scenario, the yellow areas must contain jobs $j\in S_{1}$.
+The near-equal partitioning of the rows depend on the value $m\!\!\mod 3$, which might be at times surprisingly aggravating to calculate with, which is why we have provided the reader Table 2 as an aid.
+$\left\lceil\frac{m}{3}\right\rceil$ $\left\lceil\frac{2m}{3}\right\rceil-\left\lceil\frac{m}{3}\right\rceil$ $m-\left\lceil\frac{2m}{3}\right\rceil$
+0 $\frac{m}{3}$ $\frac{m}{3}$ $\frac{m}{3}$
+1 $\frac{m+2}{3}$ $\frac{m-1}{3}$ $\frac{m-1}{3}$
+2 $\frac{m+1}{3}$ $\frac{m+1}{3}$ $\frac{m-2}{3}$
+Table 2: Numbers of rows in each partition, depending on $m\!\!\mod 3$.
+Case 1: There is no job $j\in S_{1}\cap S_{2}\cap S_{3}$ in the $m^{\prime}$-th machine.
+Let $k\in\{1,2,3\}$ be the scenario in the $m^{\prime}$-th machine attaining the makespan of $3$.
+We use Table 2 to find a lower bound.
+To be more precise, depending on $m\!\!\mod 3$, numbers from one of the three rows in this table are used to bound $S_{k}$ from below; it is at least two times one entry plus another entry (plus $3$ which comes from the $m^{\prime}$-th row).
+Therefore we have
+$p(S_{k})\geq\min\left\{\frac{2m}{3}+\frac{m}{3}+3,\frac{2m-2}{3}+\frac{m-1}{3}+3,\frac{2m-4}{3}+\frac{m+1}{3}+3\right\}>m.$
+For the next cases, we observe the following:
+Observation 6.3.
+Fix a column $k$. Then in this column, the counting index of a job $j\in S_{1}\cap S_{2}\cap S_{3}$ is always strictly larger than that of a job $j^{\prime}\notin S_{1}\cap S_{2}\cap S_{3}$.
+Case 2: There are at least two jobs $j,j^{\prime}\in S_{1}\cap S_{2}\cap S_{3}$ in the $m^{\prime}$-th machine.
+If there are at least two jobs $j,j^{\prime}\in J_{m^{\prime}}\cap S_{1}\cap S_{2}\cap S_{3}$, then one of them must have counting index larger than $\frac{m-2}{3}$ by Table 2, therefore all three columns have at least $\frac{m-2}{3}$ jobs $j\in S_{1}\cap S_{2}\cap S_{3}$ not counting the $m^{\prime}$-th row.
+In total, it must hold that
+$S_{k}\geq 3\cdot\frac{m-2}{3}+3>m.$
+Case 3: There is exactly one job $j\in S_{1}\cap S_{2}\cap S_{3}$ in the $m^{\prime}$-th machine.
+If the job $j$ has counting index $s$ larger than $\frac{m-2}{3}$ (from bottom to top), then the case goes analogous to Case 2.
+Otherwise, we branch into two subcases:
+Subcase 3.1: $j$ is not in the leftmost column. Then the other two jobs not being in $S_{1}\cap S_{2}\cap S_{3}$ implies that there are at least
+$\min\left\{\left\lceil\frac{m}{3}\right\rceil+2\left(\left\lceil\frac{2m}{3}\right\rceil-\left\lceil\frac{m}{3}\right\rceil-s\right),m-\left\lceil\frac{2m}{3}\right\rceil+2\left(\left\lceil\frac{m}{3}\right\rceil-s\right)\right\}\geq m-2s,$
+jobs that precede the $m^{\prime}$-th row in the counting order in the other two columns; the inequality can be seen via Table 2.
+Together with the three jobs in $J_{m^{\prime}}$ and the 3(s1)3(s-1) jobs that precede $j$ in the backwards counting order, we obtain
+$p(S_{k})\geq m-2s+3+3s-3>m,$
+as desired.
+Subcase 3.2: $j$ is in the leftmost column. Then the job $j^{\prime}\in J_{m^{\prime}}$ in the second column is preceded by
+$\left\lceil\frac{2m}{3}\right\rceil-\left\lceil\frac{m}{3}\right\rceil+\left(m-\left\lceil\frac{2m}{3}\right\rceil-s\right)=m-\left\lceil\frac{m}{3}\right\rceil-s$
+jobs in the second column, and the job $j^{\prime\prime}\in J_{m^{\prime}}$ in the third column is preceded by $m-\left\lceil\frac{2m}{3}\right\rceil-s$ jobs in the third column, all belonging to $S_{k}$.
+Moreover, there are at least 3s13s-1 jobs in $S_{1}\cap S_{2}\cap S_{3}$ preceding $j$ in the backwards counting order.
+Here, we used the fact that tiebreaks were broken in favor of the rightmost slots.
+Together with the three jobs in $J_{m^{\prime}}$, we can bound
+$p(S_{k})\geq 2m-\left\lceil\frac{2m}{3}\right\rceil-\left\lceil\frac{m}{3}\right\rceil-2s+3s+3>m.$
+Multiple bingo cards.
+Now we consider the general case where we have occupied multiple bingo cards. Recall that in the algorithm, the job is placed to the lowest-index bingo card where a vacant placeholder exists.
+We call a column (consisting of triangles) of a bingo card column bingo if there exists a scenario that is represented in every entry of this column.
+Then, by the minimality of the card index that we select, there can be only one non-empty and non-bingo copy of each of the three columns.
+The bottleneck is achieved when there are no bingo columns because if we drop a bingo column, the makespan decreases by at most $1$ (exactly $1$ if and only if a scenario that achieves the bingo also achieves the makespan), while average load ratio output divided by average load becomes strictly larger.
+Here, notice that in our analysis, we always constructed a lower bound using the scenario that attains makespan.
+Therefore, there is a worst-case instance with only one card.
 7 Conclusion and Outlook
@@
 Surely, the fact that the subsets of $\{1,2,3\}$ are not laminar, has been a deciding factor in our simple non-competitiveness result, as has been in several of the well-known results.
-Interestingly, restricting to the proxy competitive ratio as well as fixing the method of assigning double-scenario jobs essentially reduced the existence problem to the description of a polyhedron.
-We believe that similar techniques might prove useful for other cases and related problems as well.
+It is worth mentioning that restricting to the proxy competitive ratio as well as fixing the method of assigning double-scenario jobs essentially reduced the existence problem to the description of a polyhedron.
+We believe that this technique, if understood better, might prove useful for other cases and related problems as well.
 We have further compared the behavior of competitiveness for increasing number $K$ of scenarios and increasing number $m$ of machines.
@@
 [migration].
-In this model, we are allowed to revoke a bounded size of decisions in hindsight.
+In this model, we are allowed to revoke a bounded number of decisions in hindsight.
 Similarly, instead of minimizing the worst-case scenario, one could look into minimizing the average scenario or maximum regret.